### 09 Dec 2023 v1

# DPoser: Diffusion Model as Robust 3D Human Pose Prior

Abstract
Modeling human pose is a cornerstone in applications from human-robot interaction to augmented reality, yet crafting a robust human pose prior remains a challenge due to biomechanical constraints and diverse human movements. Traditional priors like VAEs and NDFs often fall short in realism and generalization, especially in extreme conditions such as unseen noisy poses. To address these issues, we introduce DPoser, a robust and versatile human pose prior built upon diffusion models. Designed with optimization frameworks, DPoser seamlessly integrates into various pose-centric applications, including human mesh recovery, pose completion, and motion denoising. Specifically, by formulating these tasks as inverse problems, we employ variational diffusion sampling for efficient solving. Furthermore, acknowledging the disparity between the articulated poses we focus on and structured images in previous research, we propose a truncated timestep scheduling to boost performance on downstream tasks. Our exhaustive experiments demonstrate DPoser’s superiority over existing state-of-the-art pose priors across multiple tasks. Corresponding authors: Yulun Zhang and Haoqian Wang.
Figure 1: An overview of DPoser’s versatility and performance across multiple pose-related tasks. Built on diffusion models, DPoser serves as a robust and adaptable pose prior. Shown are applications in (a) pose generation, (b) human mesh recovery, (c) motion denoising, and (d) pose completion. DPoser consistently outstrips existing priors like VPoser [35] in performance benchmarks.
1 Introduction
The accurate modeling of human pose is an essential building block in a variety of applications, ranging from human-robot interaction to augmented and virtual reality experiences. Many real-world applications rely on a prior distribution of valid human poses to perform tasks like body model fitting, motion capture, and gesture recognition. Existing methods use Gaussian Mixture Models (GMMs) [34], Variational Autoencoders (VAEs) [35], or Neural Distance Fields (NDFs) [36] to model the complex human pose prior. However, these models have limitations. GMMs can generate implausible poses due to their unbounded nature, VAEs are constrained by Gaussian assumptions, and NDFs exhibit limited generalization capabilities, requiring additional training strategies for extreme poses. These shortcomings underscore the need for a more robust and reliable human pose prior, a gap that this work aims to fill.
Diffusion models [10, 8, 12, 14] have recently gained traction for their prowess in capturing complex, high-dimensional data distributions and enabling versatile sampling techniques. Although they have been employed in specialized applications like generating realistic human motion sequences [45, 46] or serving as multi-hypothesis pose estimators based on 2D inputs [41, 43], their utility has generally been constrained. These models are designed for generation or tailored to work with specific types of input data, which limits their applicability in broader contexts. Thus far, the potential of diffusion models as a universal human pose prior remains largely untapped, and effective optimization methods for diverse applications remain unanswered.
In this work, we propose DPoser, a novel approach that leverages time-dependent denoiser learned from expansive motion capture datasets to construct a robust human pose prior. Specifically, by viewing various pose-centric tasks as inverse problems, we propose to employ variational diffusion sampling techniques [18] to incorporate DPoser as a regularization term within the existing optimization framework e.g., SMPLify [34]. Furthermore, our experimental investigations into the generative aspects of articulated human poses reveal that the major information during diffusion is concentrated on the “clean” end of the perturbed data, as opposed to the entire noise-to-clean trajectory. Motivated by this insight, we introduce a truncated annealed timestep scheduling for optimization, which demonstrates superior efficiency over conventional random and full annealed schedules across multiple downstream tasks.
In summary, our main contributions of this work can be summarized as follows:
  • We introduce DPoser, a novel framework based on diffusion models to craft a robust and flexible human pose prior, geared for seamless integration across diverse pose-related applications via test-time optimization.
  • Utilizing a fresh perspective on inverse problems, we employ variational diffusion sampling during test-time optimization and introduce truncated timestep scheduling, showing superior efficiency over traditional methods.
  • Through extensive experiments and ablation studies, we establish that DPoser not only outshines state-of-the-art (SOTA) pose priors in a variety of applications but also provides valuable insights into training techniques such as suitable rotation representations for peak performance.
2 Related works
2.1 Human Pose Priors
Human body models such as SMPL [33] serve as powerful tools for parameterizing both pose and shape, thereby offering a comprehensive framework for describing a wide array of human gestures. Within the SMPL model, body poses are captured using rotation matrices or joint angles linked to a kinematic skeleton. Adjusting these parameters enables the representation of a diverse range of human actions. However, due to inherent biomechanical constraints, plausible poses inherently reside in a lower-dimensional manifold.
Numerous approaches [34, 35, 36] have been proposed to model this intrinsic pose prior. Classical generative models like GMMs, VAEs [16], and GANs [17] have been employed to capture the complex underlying distribution, demonstrating significant improvements in downstream tasks such as human mesh recovery [49, 50]. Additionally, some research has explored task-specific conditional pose priors, augmenting the model with additional inputs like image features [42, 48], 2D joint coordinates [43], or preceding poses in motion sequences [40, 39]. Our work adopts the unconditional pose prior framework. We train DPoser on large-scale motion capture datasets without auxiliary inputs like paired images or captions, to ensure its applicability across a broad spectrum of pose-related tasks.
2.2 Diffusion Models for Pose-centric Tasks
Diffusion models [6, 8, 10, 11] have emerged as powerful tools for capturing intricate data distributions, aligning particularly well with the demands of multi-hypothesis estimation in ambiguous human poses. Notable works include DiffPose [41], which leverages a Gaussian Mixture Model-guided forward diffusion process [58] and employs a Graph Convolutional Network (GCN) [57] architecture conditioned on 2D pose sequences for 3D pose estimation by learned reverse process (i.e., generation). In a similar vein, DiffusionPose [42] and GFPose [43] employ the generation-based pipeline but take different approaches in conditioning. ZeDO [44] focuses on 2D-to-3D pose lifting, while works like Diff-HMR [48] and DiffHand [47] extend to estimating SMPL parameters and hand mesh vertices. BUDDI [62] stands out for using diffusion models to capture the joint distribution of interacting individuals and leveraging SDS loss [27, 28] for test-time optimization.
Our work, DPoser, parallels BUDDI in similar optimization implementation. However, we distinguish ourselves by introducing a wider perspective of inverse problems and proposing advanced timestep scheduling based on characteristics of human poses. Moreover, DPoser is designed to work as a flexible regularization term, making it adaptable to diverse pose-centric tasks. It should also be noted that unlike most works [44, 42, 43, 41] that focus on joint positions, DPoser targets the more complex issue of joint rotations, adding an extra layer of challenge due to the intricacy of rotation representations.
3 Preliminary: Score-based Diffusion Models
Diffusion models [3, 6, 8, 10] operationalize generative processes by inverting a predefined forward diffusion process, typically formulated as a linear stochastic differential equation (SDE). Formally, the data trajectory $\left\{\mathbf{x}(t)\in\mathbb{R}^{n}\right\}_{t\in[0,1]}$ follows the forward SDE given by:
$\mathrm{d}\mathbf{x}=\mu(t)\mathbf{x}\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}.$ (1)
Here, $\mu(t)\mathbf{x}\in\mathbb{R}^{n}$ and $g(t)\in\mathbb{R}$ represent the drift and diffusion coefficients, while $\mathbf{w}$ is a standard Wiener process.
The affine drift coefficients ensure analytically tractable Gaussian perturbation kernels, denoted by $p_{0t}(\mathbf{x}_{t}\mid\mathbf{x})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x},\sigma_{t}^{2}\mathbf{I})$, where the exact coefficients $\alpha_{t},\sigma_{t}$ can be obtained with standard techniques [7]. Using appropriately designed $\alpha_{t}$ and $\sigma_{t}$, this allows the data distribution $\mathbf{x}_{0}\sim p_{data}$ to morph into a tractable isotropic Gaussian distribution $\mathbf{x}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ via forward diffusion.
To recover data distribution $p_{data}$ from the Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$, we can simulate the corresponding reverse SDE of Eq. (1[8]:
$\mathrm{d}\mathbf{x}=[\mu(t)\mathbf{x}-g(t)^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\right)]\mathrm{d}t+g(t)\mathrm{d}\bar{\mathbf{w}}.$ (2)
The so-called score function [4], $\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\right)$, serves as an unknown term in Eq. (2) and can be approximated by a neural network parameterized as $\epsilon_{\theta}(\mathbf{x}_{t};t)\approx-\sigma_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\right)$ [^1] This parameterization is obtained from the deep connection between the noise prediction in diffusion models and score function estimation in score-based models. We provide a brief recap in the Appendix.. To learn the score functions, employing denoising score matching techniques [2], we first introduce noise to the data points as per:
$\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\epsilon,\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (3)
Subsequently, we train the time-dependent noise predictor $\epsilon_{\theta}(\mathbf{x}_{t};t)$ using an L2-loss defined as [10]:
$\mathbb{E}_{\mathbf{x}_{0}\sim p_{\mathrm{data}},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[w(t)||\epsilon-\epsilon_{\theta}(\mathbf{x}_{t};t)||_{2}^{2}\right],$ (4)
where $w(t)$ denotes a positive weighting function. Upon successful training, the score functions can be estimated and used to solve the reverse SDE (Eq. (2)). Through techniques like Euler-Maruyama discretization, we can generate novel samples by simulating the reverse SDE.
4 Methods
4.1 Learning Pose Prior with Diffusion Models
We propose to model pose prior based on the SMPL body model [33], which can be viewed as a differentiable function $[J,V]=M(\theta,\beta)$ that maps body joint angles $\theta\in\mathbb{R}^{3\times 21}$ and shape parameters $\beta\in\mathbb{R}^{10}$ to mesh vertices $V\in\mathbb{R}^{3\times 6890}$ and joint positions $J\in\mathbb{R}^{3\times 22}$. Our target is to model the distribution of joint angles $p(\theta)$. To achieve this, we train DPoser on mocap datasets, employing sub-VP SDEs introduced in ScoreSDE [8], defined as:
$\mathrm{d}\mathbf{x}=-\frac{1}{2}\beta(t)\mathbf{x}\mathrm{d}t+\sqrt{\beta(t)(1-e^{-2\int_{0}^{t}\beta(s)\mathrm{d}s})}\mathrm{d}\mathbf{w},$ (5)
where $\beta(t)$ denotes linear scheduled noise scales. The coefficients needed in Eq. (3) can be obtained as $\alpha_{t}=e^{-\frac{1}{2}\int_{0}^{t}\beta(s)\mathrm{d}s},\sigma_{t}=[1-e^{-\int_{0}^{t}\beta(s)\mathrm{d}s}]^{2}$. With the constructed forward process (Eq. (3)) to sample $\mathbf{x}_{t}$, we then apply the objective in Eq. (4) to train the noise predictor $\epsilon_{\theta}(\mathbf{x}_{t};t)$ with weights $w(t)=\sigma_{t}^{2}$ as suggested in ScoreSDE [8].
In our experiments, we find the selection of rotation representations, such as axis-angle and 6D rotations [59], to be crucial for optimal performance when coupled with suitable normalization. Despite experimenting with incorporating body models via auxiliary losses, as in prior work [35, 39], we note that it imposes a computational burden without evident gains in downstream task performance. Further methodological nuances are detailed in the Appendix.
4.2 Optimization Leveraging Diffusion Priors
Figure 2: Overview of the DPoser Methodology. Panel (a) presents three tasks: human mesh recovery, pose completion, and motion denoising, with omissions like camera optimization and bending loss for clarity. Panel (b) demonstrates the DPoser regularization process, introducing noise to the current pose and applying a one-step denoiser to achieve a denoised pose. $L_{\text{DPoser}}$ is computed between the denoised and current pose. Panel (c) outlines the optimization process from initial to current and fitted poses via loss minimization.
The acquired score functions or noise predictors, denoted as $\epsilon_{\theta}(\mathbf{x}_{t};t)$, permit the direct generation of plausible poses through Eq. (2). Yet, the broader integration of diffusion priors into general optimization frameworks remains an open avenue. We address this by reframing pose-related tasks as inverse problems and applying variational diffusion sampling techniques [18] for efficient resolution.
Inverse problem formulation. Consider an original signal $\mathbf{x}_{0}$. Inverse problems can be encapsulated by Eq. (6) as:
$\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})+\mathbf{n},\quad\mathbf{y},\mathbf{n}\in\mathbb{R}^{d},~{}\mathbf{x}_{0}\in\mathbb{R}^{n},$ (6)
where $\mathcal{A}$ symbolizes the measurement operator and $\mathbf{n}$ constitutes noise, assumed to be white Gaussian $\mathcal{N}(\mathbf{0},\sigma_{n}^{2}\mathbf{I})$ in our context. As an illustration, inverse kinematics can be considered as inferring joint rotations (or body poses in SMPL [33]) from observed joint positions. Moreover, the 2D-3D lifting task can be modeled as a linear inverse problem, with $\mathcal{A}$ representing a camera projection model. To recover the original signal $\mathbf{x}_{0}$, in the Bayesian framework, our target is to sample from the posterior distribution $p\left(\mathbf{x}_{0}\mid\mathbf{y}\right)$.
Diffusion-based inverse problem solving. Various techniques [26, 21, 19, 20, 22, 18] have been explored to simulate this posterior sampling process based on diffusion priors. For some noiseless linear inverse problems, such as completion, alternating projections onto measurement subspaces are commonly employed [8, 21, 20]. For instance, in hand pose completion, a standard iterative generation process using diffusion models can be applied, replacing known parts of the pose at each iteration.
However, such projections are inconvenient to build general optimization frameworks concerned in our work. To navigate these challenges, we adopt variational diffusion sampling [18], efficient in computation and straightforward in implementation. Specifically, it employs a variational distribution $q\left(\mathbf{x}_{0}\mid\mathbf{y}\right):=\mathcal{N}(\mu,\sigma^{2}\mathbf{I})$ and aims to minimize the Kullback-Leibler (KL) divergence between this variational distribution and the true posterior, mathematically expressed as $KL\big{(}q\left(\mathbf{x}_{0}\mid\mathbf{y}\right)\parallel p\left(\mathbf{x}_{0}\mid\mathbf{y}\right)\big{)}$.
Informed by existing theorems concerning the likelihood properties of diffusion models [9], and under the assumption of zero variance ($\sigma\approx 0$), the optimization problem of seeking $\mathbf{x}_{0}$ (i.e., $\mu$) can be formulated as minimizing [18]:
$\|\mathbf{y}-\mathcal{A}(\mathbf{x}_{0})\|^{2}+w_{t}(\mathtt{sg}[\epsilon_{\theta}(\mathbf{x}_{t};t)-\epsilon])^{\top}\mathbf{x}_{0},$ (7)
where $w_{t}$ denotes the loss weights and $\epsilon$ is sampled from the standard Gaussian distribution. Here, $\mathtt{sg}$ signifies the stopped-gradient operator, indicating that backpropagation through the trained diffusion models is not required. The optimization procedure initiates by selecting a timestep $t$ and applying a perturbation to the target $\mathbf{x}_{0}$ as per Eq. (3), resulting in $\mathbf{x}_{t}$. Subsequently, the gradients $[\epsilon_{\theta}(\mathbf{x}_{t};t)-\epsilon]$ are applied to the optimization variable $\mathbf{x}_{0}$. In a nutshell, this framework [18] provides a flexible yet robust strategy for employing diffusion priors in generic optimization problems, serving as a cornerstone for our work.
DPoser regularization. To shed more light on the working mechanism, we reformulate the regularization term as [^2] We detail another perspective of Score Distillation Sampling (SDS) [27, 28] to understand our DPoser regularization in the Appendix.:
$L_{\mathrm{DPoser}}$ $=w_{t}||\mathbf{x}_{0}-\mathtt{sg}[\mathbf{\hat{x}}_{0}(t)]||_{2}^{2},\text{where}$ (8)
$\mathbf{\hat{x}}_{0}(t)$ $=\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\theta}(\mathbf{x}_{t};t)}{\alpha_{t}}.$ (9)
Here, $\mathbf{\hat{x}}_{0}(t)$ serves as a one-step denoised estimate, leveraging the trained diffusion model $\epsilon_{\theta}(\mathbf{x}_{t};t)$. We show that the gradient of our DPoser regularization aligns with Eq. (7).
Proof: Differentiating Eq. (8) with respect to $\mathbf{x}_{0}$ yields:
$\nabla_{\mathbf{x}_{0}}L_{\mathrm{DPoser}}$ $=2w_{t}(\mathbf{x}_{0}-\mathbf{\hat{x}}_{0}(t))$
$=2w_{t}(\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon}{\alpha_{t}}-\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\theta}(\mathbf{x}_{t};t)}{\alpha_{t}})$
$=2w_{t}\frac{\sigma_{t}}{\alpha_{t}}(\epsilon_{\theta}(\mathbf{x}_{t};t)-\epsilon)$
$\propto(\epsilon_{\theta}(\mathbf{x}_{t};t)-\epsilon).$ (10)
Therefore, our DPoser regularization $L_{\mathrm{DPoser}}$ can be implemented as an L2-loss. In this view, during optimization, DPoser works by pulling the current pose $\mathbf{x}_{0}$ closer to the learned plausible pose distribution through denoising.
DPoser across pose-related tasks. Capitalizing on DPoser’s inherent flexibility, we adapt it to various human pose tasks, with a spotlight on human mesh recovery as illustrated in Fig. 2. For an exhaustive examination of DPoser’s utility across different pose-related applications including pose completion and motion denoising, we direct the reader to our Appendix.
By adapting the objective function from the SMPLify framework [34], we supplant the originally used GMM prior with our own DPoser regularization term, $L_{\mathrm{DPoser}}$, while also forgoing the intricate interpenetration error term. The optimization objective, incorporating both pose $\theta$ and shape $\beta$ parameters in SMPL [33], is articulated as:
$L(\theta,\beta)=L_{J}+w_{\theta}L_{\theta}+w_{\beta}L_{\beta}+w_{\alpha}L_{\mathrm{DPoser}}.$ (11)
Herein, the reprojection loss $L_{J}$ serves as the data fidelity term, mathematically expressed as:
$L_{J}$ $=\sum_{i\in\text{Joints}}\lambda_{i}\rho\left(\Pi_{C}\left(M_{J}(\theta,\beta)_{i}\right)-J^{\text{est}}_{i}\right),$ (12)
where $M_{J}(\theta,\beta)$ and $\Pi_{C}$ symbolize the 3D joint positions regressed from SMPL and the 3D-to-2D camera projection function, respectively. $J^{\text{est}}$ are the 2D keypoints estimated using an off-the-shelf 2D pose estimator (OpenPose [60] in our experiments), and $\lambda_{i}$ indicates the confidence score of joint $i$. The Geman-McClure error function [61] $\rho$ serves to robustly quantify discrepancies in 2D joint positions.
The bending term $L_{\theta}$ is incorporated to penalize excessive bending at the elbows and knees, formulated as $L_{\theta}=\sum_{i\in\text{(elbows, knees)}}\exp(\boldsymbol{\theta}_{i})$. The shape regularizer $L_{\beta}$ is a simple L2 norm, i.e., $L_{\beta}=\|\beta\|_{2}^{2}$. The constant weights for prior terms are denoted as $w_{\theta},w_{\beta}$ and $w_{\alpha}$, respectively.
Contrasting SMPLify’s GMM-based prior, we employ $L_{\mathrm{DPoser}}$ to draw the pose closer to our learned manifold. Given the structure of $L_{\mathrm{DPoser}}$ (as seen in Eq. (8)), a crucial aspect lies in judiciously selecting the diffusion timestep $t$ during the iterative optimization process. We address this concern by introducing our novel truncated timestep scheduling strategy in the subsequent section.
4.3 Test-time Truncated Timestep Scheduling
Figure 3: Illustration of the rationale behind our proposed truncated timestep scheduling. We employ the deterministic DDIM sampler [11] with limited steps and assess the quality of generated poses using the Self-Intersection percentage (SI).
Algorithm 1 Test-time Optimization with DPoser
1:A trained diffusion model $\epsilon_{\theta}(\mathbf{x}_{t};t)$, task-specific loss $L_{\text{task}}$, range of diffusion timesteps $[t_{\text{max}},t_{\text{min}}]$, number of optimization iterations $N$, DPoser weights $w_{t}$
2:Initialization of parameters $\mathbf{x}_{0}$
3:for $\text{iter}=0,1,\ldots,N-1$ do
4:     $t\leftarrow t_{\text{max}}-\frac{(t_{\text{max}}-t_{\text{min}})\times\text{iter}}{N-1}$triangleright Timestep scheduling
5:     Sample $\epsilon\sim\mathcal{N}(0,I)$
6:     $\mathbf{x}_{t}\leftarrow\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\epsilon$triangleright Forward diffusion
7:     $\mathbf{\hat{x}}_{0}(t)\leftarrow\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\theta}(\mathbf{x}_{t};t)}{\alpha_{t}}$triangleright One-step denoiser
8:     $L_{\text{DPoser}}\leftarrow w_{t}\lVert\mathbf{x}_{0}-\text{sg}[\mathbf{\hat{x}}_{0}(t)]\rVert_{2}^{2}$
9:     $L_{\text{total}}\leftarrow L_{\text{task}}+L_{\text{DPoser}}$
10:     Update $\mathbf{x}_{0}$ via backpropagation on $L_{\text{total}}$
11:end for
12:return $\mathbf{x}_{0}$
Insights from pose generation. In optimizing pose-related tasks, one fundamental challenge arises from the gap between the image-based data and the articulated pose parameters we manipulate. Prior studies in image domains [15] have noted that earlier timesteps (i.e., smaller $t$) are typically aligned with perceptual content, while the later timesteps are more closely associated with the refinement of details in the diffusion process. As illustrated in Fig. 3, our experimental foray into employing diffusion models for pose generation yields an interesting insight into pose domains: the concentration of meaningful information resides at the “clean” side of the diffusion trajectory, rather than being dispersed uniformly from noise to clean states. This observation serves as the catalyst for our introduction of a truncated annealed timestep scheduling mechanism.
Truncated timestep scheduling. Departing from traditional diffusion models that uniformly sample across the continuum from $1.0$ to $0.0$, we introduce a selective sampling strategy. Specifically, we truncate the timestep range to intervals such as $0.2\to 0$, thereby focusing on the data segments most laden with meaningful information. Based on the linear annealed scheduling, the truncated timestep $t$ for each optimization step can be formulated as:
$t=t_{\text{max}}-\frac{(t_{\text{max}}-t_{\text{min}})\times\text{iter}}{N-1}.$ (13)
Here, $N$ signifies the total optimization iterations, while iter represents the current optimization iteration. We summarize our DPoser optimization framework as Algorithm 1, where the truncated range $[t_{\text{max}},t_{\text{min}}]$ is typically set to $[0.2,0.001]$.
Through ablation studies, we prove that the proposed truncated timestep scheduling offers significant advantages over the random, fixed, and full annealed ones utilized in previous works [62, 18, 19, 8], delivering more efficient optimization across a wide range of downstream tasks. This is achieved by concentrating computation on the most informative portions of the data trajectory, as informed by our preliminary sampling investigations on pose generation.
5 Experiments
In this section, we first elaborate on the implementation details and evaluation metrics. We showcase the robustness and versatility of DPoser across a spectrum of pose-centric applications, including pose generation, human mesh recovery, pose completion, and motion denoising. Additionally, we conduct an ablation study to delve into the key elements contributing to our model’s performance.
Sample source APD $\uparrow$ SI $\downarrow$
Real-world (AMASS) [52] 15.44 0.79
GMM [34] 16.28 1.54
VPoser [35] 10.75 1.51
Pose-NDF [36] 18.75 1.97
DPoser (ours) 14.28 1.21
DPoser (ours)* 19.03 1.13
Table 1: Comparative analysis of pose generation metrics. The discrepancy between visual impressions and APD/SI metrics is discussed, with reference to Fig. 4. *Indicates the use of a DDIM sampler [11] with a reduced 10-step discretization.
(a) AMASS samples
(b) DPoser (ours)
(c) DPoser (ours)*
(d) GMM
(e) Pose-NDF
(f) VPoser
Figure 4: Qualitative comparison of generated human poses: (b) illustrates naturalistic poses closely aligned with real-world data, whereas (c) shows poses that, despite superior APD and SI metrics, lack natural appearance. *We use a DDIM sampler [11] with only 10 steps.
5.1 Experimental Setup
Implementation details. We train our DPoser model on the AMASS dataset [52], adhering to the same training partition as previous works [35, 36]. The model employs axis-angle representation for joint rotations, which we normalize to have zero mean and unit variance. The architecture consists of a fully connected neural network with approximately 8.28M parameters. It draws inspiration from GFPose [43] but omits conditional input pathways for our unconditional setting. To stabilize training, we use an exponential moving average with a decay factor of 0.9999, as advised by ScoreSDE [8]. The Adam optimizer, a learning rate of $2\times 10^{-4}$, and a batch size of 1280 govern the optimization process. The training of 800,000 iterations takes roughly 8 hours on a single Nvidia RTX 3090Ti GPU.
Evaluation metrics. We adopt task-specific metrics:
  • Pose Generation: Diversity and fidelity are evaluated using Average Pairwise Distance (APD) [37] and Self-Intersection rates (SI) [36], respectively.
  • Human Mesh Recovery: Metrics include Mean Per Joint Position Error (MPJPE) and Procrustes-aligned MPJPE (PA-MPJPE).
  • Pose Completion: We utilize both MPJPE and Mean Per-Vertex Position Error (MPVPE), for masked areas.
  • Motion Denoising: Metrics are MPVPE and MPJPE.
All errors are reported in millimeter units.
5.2 Pose Generation
To commence, we delve into the capabilities of our DPoser model by generating samples from the learned manifold. Employing a standard Euler-Maruyama discretization with 1000 steps, we assess both the diversity and realism of the generated poses (Fig. 4). While DPoser’s outputs are visually diverse and realistic, poses generated from competing methods like GMM [34] and Pose-NDF [36] fall short in naturalism, and VPoser [35] exhibits limited diversity.
Interestingly, quantitative metrics such as APD and SI (Tab. 1) do not always corroborate our qualitative findings. For instance, a 10-step DDIM sampler [11]—suboptimal by design—outperformed real-world data [52] in APD, which we attribute to the generation of exaggerated poses.
In summary, our findings underscore the need for a balanced evaluation strategy that merges quantitative metrics with qualitative observations. Beyond generation, DPoser also performs well in interpolation tasks, based on corresponding ODEs with marginal densities akin to Eq. (2[8]. Additional experiments are detailed in the Appendix.
Methods PA-MPJPE $\downarrow$ MPJPE $\downarrow$
GMM [34] 57.53 79.66
VPoser [35] 61.10 85.25
Pose-NDF [36] 61.67 85.09
DPoser (ours) 57.28 84.79
CLIFF [51] 57.30 92.86
+ GMM [34] 55.54 78.32
+ VPoser [35] 56.12 79.68
+ Pose-NDF [36] 55.50 78.39
+ DPoser (ours) 54.87 77.62
Table 2: Performance comparison of human mesh recovery on the EHF dataset [35] using two initialization methods.
Methods Occ. left leg Occ. legs Occ. arms Occ. trunk
MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
Pose-NDF [36] ($S=1$) 499.35 467.26 503.66 470.82 378.61 312.16 319.65 208.35
Pose-NDF [36] ($S=1$${}^{\dagger}$ 597.72 558.45 606.68 566.01 370.3 283.43 346.18 228.54
VPoser [35] ($S=1$) 465.98 437.92 464.38 436.06 387.14 293.17 298.63 195.16
VPoser [35] ($S=1$${}^{\dagger}$ 194.44 177.01 199.78 183.34 205.17 160.60 75.42 48.30
VPoser [35] ($S=5$) 458.92 431.31 456.01 427.90 374.32 283.90 288.96 188.26
VPoser [35] ($S=10$) 455.44 427.79 450.88 423.04 371.26 281.62 286.55 186.28
DPoser (ours) ($S=1$) 85.46 80.66 107.07 100.97 108.27 84.37 46.37 29.99
DPoser (ours) ($S=5$) 48.61 44.79 74.67 69.64 75.92 59.90 26.59 16.68
DPoser (ours) ($S=10$) 39.77 36.15 65.24 60.54 66.77 52.75 21.58 13.41
Table 3: Performance metrics for pose completion under varying occlusion scenarios. $S$ denotes the number of hypotheses. ${}^{\dagger}$ Comparison methods are implemented in the optimization paradigm.
5.3 Human Mesh Recovery
(a) EHF dataset
(b) MSCOCO dataset
(c) 3DPW dataset
(d) UBody dataset
Figure 5: Qualitative evaluations of HMR leveraging DPoser as pose prior. CLIFF [51] serves as the optimization initializer.
We probe the efficacy of DPoser in HMR, focusing on estimating human pose and shape from monocular images. We conduct experiments on the EHF dataset [35] and benchmark our method against existing SOTA priors. Our optimization-based framework incorporates two initialization paradigms: (1) a baseline initialization that utilizes mean pose values and a default camera setup, and (2) an advanced initialization scheme that leverages CLIFF [51], a pre-trained regression-based model tailored for HMR.
As shown in Tab. 2, DPoser demonstrates robust performance in HMR. With baseline initialization, it closely rivals the strong constraints of the GMM prior [34]. When enhanced by advanced initialization via CLIFF, DPoser surpasses established deep priors like Pose-NDF [36]. Fig. 5 visually confirms DPoser’s superior efficacy and adaptability across multiple datasets including EHF [35], MSCOCO [54], 3DPW [55], and UBody [53]. Further qualitative comparisons are available in the Appendix.
Methods AMASS [52] HPS [56]
VPoser [35] 23.42 22.78
Pose-NDF [36] 22.13 23.82
MVAE [40] 26.80 N/A
HuMoR [39] 22.69 N/A
DPoser (ours) 19.87 20.54
Table 4: Performance metrics (MPJPE) for motion denoising.
5.4 Pose Completion
Figure 6: Qualitative evaluations of pose completion. Three hypotheses are drawn from VPoser [35] and our DPoser.
Timestep scheduling HMR Pose Completion Motion Denoising
PA-MPJPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
Random 68.21 98.57 109.53 102.37 43.33 23.87
Fixed 57.72 85.11 40.44 36.70 45.69 22.54
Full annealed 61.15 85.27 42.97 39.42 39.72 20.80
Truncated annealed 57.28 84.79 39.77 36.15 38.21 19.87
Table 5: Comparative performance of different timestep scheduling across pose-related tasks.
Noise std AMASS [52] HPS [56]
20.00 31.93/13.64 31.93/13.45
40.00 63.81/19.87 63.81/20.54
100.00 159.78/33.18 159.78/35.32
Table 6: DPoser in motion denoising under varying noise scales. MPJPE is reported as before/after applying DPoser denoising.
In practical scenarios like those encountered in the UBody dataset [53] (refer to Fig. 4(d)), HMR algorithms often grapple with occlusions leading to incomplete 3D pose estimates. In this context, our ambition is to recover full 3D poses from such partially observed data, where the unseen parts are initialized with random noise.
Given that Pose-NDF [36] and VPoser [35] have not been previously scrutinized for this task, we introduce two experimental paradigms—generation-based and optimization-based. In the optimization regimen, we incorporate an L2-norm fidelity term for the observed pose and constrain the entire pose using these deep priors. In contrast, the generation-based scheme for VPoser involves encoding the whole pose, followed by decoding and sampling from the resulting Gaussian distributions. For Pose-NDF, we iteratively project the whole pose onto the learned manifold, replacing the unmasked components with known data. For DPoser, we report optimization-based outcomes while the generation-based results employing other inverse problem solvers [8, 20, 19] can be checked in the Appendix.
Given the task’s inherent ambiguity, we sample multiple solutions and select the best-fitting one, except the deterministic Pose-NDF. Tab. 3 and Fig. 6 demonstrate DPoser’s superiority in pose completion under occlusions, offering multiple plausible solutions, unlike VPoser [35] and Pose-NDF [36], with the latter facing generalization issues when applied to such unseen noisy poses.
5.5 Motion Denoising
Though not initially designed for temporal tasks, DPoser shows remarkable proficiency in motion denoising. The task aims to estimate clean body poses from noisy 3D joint positions in motion capture sequences. Adhering to the setup outlined in HuMoR [39], we utilize 60-frame sequences from the AMASS [52] test dataset and artificially introduce Gaussian noise with a standard deviation of 40 mm to the 3D joint positions. Moreover, we conduct experiments on HPS datasets [56] without additional training to validate the generalization. Our optimization loss is composed of a data fidelity term, a temporal consistency term, and our pose prior, as discussed in the Appendix.
As presented in Tab. 4, DPoser sets a new standard in motion denoising, outperforming even specialized motion priors like HuMoR [39]. To further confirm the robustness of DPoser, we conduct evaluations under varying conditions to gauge DPoser’s denoising capabilities. The results, detailed in Tab. 6, reveal that DPoser consistently achieves significant reductions in MPJPE, maintaining robust performance even under extreme noise conditions, specifically at a noise standard deviation of 100.00 mm. We direct the reader to our Appendix for qualitative assessments.
5.6 Ablation Study
In our ablation study, we initially focus on the impact of truncated timestep scheduling on DPoser’s performance. This involves contrasting our proposed scheduling strategy against three established methods—random, fixed, and full annealed scheduling [62, 18, 19, 8]. As Tab. 5 demonstrates, our strategy consistently outperforms these alternatives across all evaluated tasks. Further, we delve into the training aspects of DPoser, such as rotation representations and the integration of an auxiliary loss akin to HuMoR [39]. Additionally, using the same trained prior, we benchmark DPoser against SOTA diffusion-based inverse problem solvers [8, 20, 19] on the pose completion task. The results underscore that DPoser not only proves versatile across multiple pose-related tasks but also excels in performance. Detailed findings and analyses from these ablation studies are presented in the Appendix.
6 Conclusion
We introduce DPoser, a robust and vetsatile human pose prior framework grounded in diffusion models. Engineered for flexibility, DPoser shines across multiple pose-related tasks, enhanced by our innovative truncated timestep scheduling for test-time optimization. Comprehensive experiments substantiate DPoser’s superior performance over existing state-of-the-art pose priors, establishing it as an effective solution for a wide range of applications.
Limitation and future work. While our framework benefits from variational diffusion sampling [18], it also shares its limitations, such as the mode-seeking behavior. Future research could look into enhancing solution diversity via techniques like particle-based variational inference [30, 31]. Furthermore, within the broader context of inverse problems, a plethora of existing methods [22, 23, 25, 24] could be adapted to leverage our diffusion-based prior. Exploring these methods holds great potential for future progress.
Supplementary Material:
DPoser: Diffusion Model as Robust 3D Human Pose Prior
A Parameterization of Score-based Diffusion Models
In the seminal work by Song et al. [8], it is demonstrated that both score-based generative models [6] and diffusion probabilistic models [10] can be understood as discretized versions of stochastic differential equations (SDEs) defined by score functions. This unification allows the training objective to be interpreted either as learning a time-dependent denoiser or as learning a sequence of score functions that describe increasingly noisy versions of the data.
We begin by revisiting the training objective for score-based models [6] to elucidate the link with diffusion models [10]. Consider the transition kernel of the forward diffusion process $p_{0t}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I})$. Our goal is to learn score functions $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ through a neural network $s_{\theta}(\mathbf{x}_{t};t)$, by minimizing the L2 loss as follows (we omit the expectation operator for conciseness) :
$\mathbb{E}\left[w(t)||s_{\theta}(\mathbf{x}_{t};t)-\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\right)||_{2}^{2}\right].$ (14)
Here, $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\epsilon$, where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
Based on denoising score matching [2], we know the minimizing objective Eq. (14) is equivalent to the following tractable term:
$\mathbb{E}\left[w(t)||s_{\theta}(\mathbf{x}_{t};t)-\nabla_{\mathbf{x}_{t}}\log p_{0t}(\mathbf{x}_{t}|\mathbf{x}_{0})||_{2}^{2}\right].$ (15)
To link this with the noise predictor $\epsilon_{\theta}(\mathbf{x}_{t};t)$ in diffusion models, we can employ the reparameterization $s_{\theta}(\mathbf{x}_{t};t)=-\frac{\epsilon_{\theta}(\mathbf{x}_{t};t)}{\sigma_{t}}$. Then, Eq. (15) can be simplified as follows:
$w(t)||-\frac{\epsilon_{\theta}(\mathbf{x}_{t};t)}{\sigma_{t}}-\nabla_{\mathbf{x}_{t}}\log p_{0t}(\mathbf{x}_{t}\mid\mathbf{x}_{0})||_{2}^{2}$
$=$ $w(t)||-\frac{\epsilon_{\theta}(\mathbf{x}_{t};t)}{\sigma_{t}}+\frac{(\mathbf{x}_{t}-\alpha_{t}\mathbf{x}_{0})}{\sigma_{t}^{2}}||_{2}^{2}$
$=$ $w(t)||-\frac{\epsilon_{\theta}(\mathbf{x}_{t};t)}{\sigma_{t}}+\frac{\sigma_{t}\epsilon}{\sigma_{t}^{2}})||_{2}^{2}$
$=$ $\frac{w(t)}{\sigma_{t}^{2}}||\epsilon_{\theta}(\mathbf{x}_{t};t)-\epsilon)||_{2}^{2}$ (16)
The resulting form of Eq. (16) aligns precisely with the noise prediction form of diffusion models [10] (refer to Eq. (4)). This implies that by training $\epsilon_{\theta}(\mathbf{x}_{t};t)$ in a diffusion model context, we simultaneously get a handle on the score function, approximated as $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\approx-\frac{\epsilon_{\theta}(\mathbf{x}_{t};t)}{\sigma_{t}}$.
B View DPoser as Score Distillation Sampling
Strategy HMR Pose Completion Motion Denoising
PA-MPJPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
1step 57.28 84.79 39.77 36.15 38.21 19.87
5steps 57.61 85.14 41.30 37.58 40.22 21.21
10steps 57.66 85.07 41.28 37.39 40.69 21.34
Table S-1: Efficacy of different denoising steps in DPoser’s optimization.
Interestingly, the gradient of DPoser in Eq. (10) coincides with Score Distillation Sampling (SDS) [27, 28], which can be interpreted as aiming to minimize the following KL divergence:
$KL\big{(}p_{0t}\left(\mathbf{x}_{t}\mid\mathbf{x}_{0}\right)\parallel p_{t}^{\mathtt{SDE}}\left(\mathbf{x}_{t};\theta\right)\big{)},$ (17)
where $p_{t}^{\mathtt{SDE}}\left(\mathbf{x}_{t};\theta\right)$ denote the marginal distribution whose score function is estimated by $\epsilon_{\theta}(\mathbf{x}_{t};t)$. For the specific case where $t\to 0$, this term encourages the Dirac distribution $\delta(\mathbf{x}_{0})$ (i.e., the optimized variable) to gravitate toward the learned data distribution $p_{0}^{\mathtt{SDE}}\left(\mathbf{x}_{0};\theta\right)$, while the Gaussian perturbation like Eq. (17) softens the constraint. Building on this understanding, we can borrow advanced techniques from SDS [27, 28]—a rapidly evolving area ripe for methodological innovations [31, 32, 29]. To extend this, we experiment with a multi-step denoising strategy adapted from HiFA [29], substituting our original one-step denoising process. This alternative, however, yields suboptimal results across all evaluation metrics, as demonstrated in Tab. S-1. A plausible explanation could be that our proposed truncated timestep scheduling effectively manages low noise levels (i.e., small $t$), thus negating the need for more denoising steps.
C Experimental Details
This section elaborates on the specifics of our pose completion and motion denoising experiments, and also highlights DPoser’s proficiency in pose interpolation tasks.
C.1 Pose completion
For partial observations $\mathbf{y}$, the measurement operator $\mathcal{A}$ is modeled as a mask matrix $M\in\mathbb{R}^{d\times n}$. Based on our optimization framework (i.e., Algorithm 1), we define the task-specific loss, $L_{\text{comp}}$, as follows:
$L_{\text{comp}}=||M\mathbf{x}_{0}-\mathbf{y}||_{2}^{2}.$ (18)
Here, $\mathbf{x}_{0}$ denotes the whole body pose $\theta$ we try to recover, where the unseen parts are initialized as random noise. In ourthe following ablated studies, if not specified, the evaluation is performed using 10 hypotheses on the AMASS [52] test set with left leg occlusion.
C.2 Motion denoising
Methods AMASS [52] HPS [56]
20mm 100mm 20mm 100mm
VPoser [35] 15.20 49.10 17.24 46.69
Pose-NDF [36] 13.84 46.10 15.62 47.50
DPoser (ours) 13.64 33.18 13.45 35.32
Table S-2: Performance comparison of motion denoising under varying noise scales. MPJPE is reported afters denoising.
Adhering to Pose-NDF settings [36], we aim to refine noisy joint positions $J_{\text{obs}}^{t}$ over $N$ frames to obtain clean poses $\theta^{t}$, initialized from mean poses in SMPL with small noise. We formulate the task-specific loss combining an observation fidelity term $L_{\text{obs}}$ and a temporal consistency term $L_{\text{temp}}$:
$L_{\text{obs}}$ $=\sum_{t=0}^{N-1}||M_{J}(\theta^{t},\beta_{0})-J_{\text{obs}}^{t}||_{2}^{2},$ (19)
$L_{\text{temp}}$ $=\sum_{t=1}^{N-1}||M_{J}(\theta^{t-1},\beta_{0})-M_{J}(\theta^{t},\beta_{0})||_{2}^{2},$ (20)
where $M_{J}$ denotes the 3D joint positions regressed from SMPL [33] and $\beta_{0}$ is the constant mean shape parameters.
In complement to the comparative analysis presented in Tab. 4, we extend our evaluation to include scenarios with varying noise levels. This extended examination, detailed in Tab. S-2, showcases DPoser’s exceptional performance against state-of-the-art (SOTA) pose priors, especially under conditions of high noise, manifesting DPoser’s resilience to noise.
C.3 Pose interpolation
(a) Pose reconstruction.
(b) Interpolation between two poses.
Figure S-1: Visual results of pose reconstruction and interpolation.
For the forward SDE (Eq. (1)), there exists a probability flow ODE with the same marginal probability densities corresponding to it, expressed as: [8]
$\mathrm{d}\mathbf{x}=[\mu(t)\mathbf{x}-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\right)]\mathrm{d}t.$ (21)
By replacing $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ with our approximation $-\frac{\epsilon_{\theta}(\mathbf{x}_{t};t)}{\sigma_{t}}$, this ODE can be treated as a neural ODE [5], which is fully invertible. By integrating through it, we can encode any data point $\mathbf{x}_{0}$ into its latent representation $\bar{\mathbf{x}}_{1}$ and decode it through its reverse-time counterpart.
For interpolation, we start with two poses, $\mathbf{x}_{0}^{0}$ and $\mathbf{x}_{0}^{1}$, and encode them into latent representations $\bar{\mathbf{x}}_{1}^{0}$ and $\bar{\mathbf{x}}_{1}^{1}$. Utilizing spherical linear interpolation (Slerp), we obtain interpolated latents $\bar{\mathbf{x}}_{1}^{t}=\text{Slerp}(\bar{\mathbf{x}}_{1}^{0},\bar{\mathbf{x}}_{1}^{1};t)$ for $t\in[0,1]$, which are then decoded to produce novel poses. The Slerp function is defined as:
$\text{Slerp}(A,B;t)=\frac{\sin((1-t)\omega)}{\sin\omega}A+\frac{\sin(t\omega)}{\sin\omega}B,$ (22)
where $\omega=\arccos(A\cdot B/\|A\|\|B\|)$.
To validate the informativeness of our encoding, we employ MPJPE and MPVPE metrics for reconstruction evaluation. Compared to VPoser [35], a VAE-based prior, DPoser exhibits superior performance with 4.79mm and 6.58mm errors, against 9.85mm and 12.16mm, respectively. Qualitative results further confirm DPoser’s strong reconstruction and smooth interpolation capabilities, as shown in Fig. S-1.
D Ablated DPoser Training
Normalization HMR Pose Completion Motion Denoising
PA-MPJPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
w/o norm 58.62 83.75 60.70 55.47 44.82 24.04
min-max 59.55 85.63 58.18 53.06 42.70 21.29
z-score 57.66 85.47 52.86 48.27 38.57 20.24
Table S-3: Evaluation of DPoser’s performance under different normalization methods, specifically for the axis-angle rotation representation.
Representation HMR Pose Completion Motion Denoising
PA-MPJPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
axis-angle 57.28 84.79 52.14 47.60 38.21 19.87
6D rotations 58.02 85.93 56.28 51.54 38.44 20.12
Table S-4: Comparative performance of rotation representations under z-score normalization across multiple tasks and metrics.
Figure S-2: Evolution of MPJPE during DPoser training, evaluated on pose completion. The evaluation is conducted on the AMASS [52] test set, utilizing 10 hypotheses and simulating occlusion scenarios affecting both legs.
Building on the insights shared in Sec. 4.1 in our main text, this part dissects the impact of different rotation representations and normalization techniques on DPoser’s performance. Initially, we examine axis-angle representation, comparing various normalization strategies: min-max scaling, z-score normalization, and no normalization. Our findings, summarized in Tab. S-3, indicate that z-score normalization is generally the most effective. Subsequently, using this optimal normalization, we explore 6D rotations [59] as an alternative. As evidenced by Tab. S-4, axis-angle representation offers superior performance. This preference can be attributed to the effective modeling capabilities of diffusion models, along with the inherent advantages of axis-angle in capturing bounded joint rotations.
Inspired by HuMoR [39], we experiment with integrating the SMPL body model [33] as a regularization term during DPoser’s training. Alongside the prediction of additive noise, as outlined in Eq (4), we employ a 10-step DDIM sampler [11] to recover a “clean” version of the pose, denoted as $\tilde{\mathbf{x}}_{0}$, from the diffused $\mathbf{x}_{t}$. The regularization loss aims to minimize the discrepancy between the original and recovered poses under the SMPL body model $M$:
$L_{\mathrm{reg}}=$ $||M_{J}(\tilde{\mathbf{x}}_{0},\beta_{0})-M_{J}(\mathbf{x}_{0},\beta_{0})||_{2}^{2}$
$+||M_{V}(\tilde{\mathbf{x}}_{0},\beta_{0})-M_{V}(\mathbf{x}_{0},\beta_{0})||_{2}^{2}.$
Here, $\beta_{0}$ represents the mean shape parameters in SMPL. To account for denoising errors, we scale the regularization loss by $\mathrm{log}(1+\frac{\alpha_{t}}{\sigma_{t}})$, thereby increasing the weight for samples with smaller $t$ values (less noise). Fig. S-2 visualizes the impact of this regularization on MPJPE during the training, specifically for pose completion tasks with occlusion of both legs. We observe that weighted regularization offers slight performance gains in the early training process, while the absence of weighting introduces instability and deterioration in results. Despite these insights, the computational cost of incorporating the SMPL model—especially for our large batch size of 1280—makes the training approximately 8 times slower. Therefore, we opted not to include this regularization in our main experiments.
E Extended DPoser Optimization
Methods Occ. left leg Occ. legs Occ. arms Occ. trunk
MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
ScoreSDE [8] 52.86 48.27 80.68 74.74 82.38 65.81 27.97 17.27
DPS [19] 41.99 38.08 66.16 61.28 64.75 51.75 25.68 15.90
MCG [20] 51.97 47.46 80.07 74.14 82.22 65.58 27.93 17.27
DPoser (ours) 39.77 36.15 65.24 60.54 66.77 52.75 21.58 13.41
Table S-5: Comparative evaluation of diffusion-based solvers for pose completion.
In addressing pose-centric tasks as inverse problems, we propose a versatile optimization framework, which employs variational diffusion sampling as its foundational approach [18]. Our exploration extends to an array of diffusion-based methodologies for solving these complex inverse problems. Among the techniques considered are ScoreSDE [8], MCG [20], and DPS [19]. These methods augment standard generative processes with observational data, either by employing gradient-based guidance or back-projection techniques. We compare these methods with our DPoser for pose completion tasks. Our findings, captured in Tab. S-5, reveal that DPoser outperforms the competitors under most occlusion conditions. Consequently, DPoser emerges not merely as a universally applicable solution to pose-related tasks, but also as an exceptionally efficient one.
It is worth mentioning that methods rooted in generative frameworks [8, 20, 19, 21] can pose challenges for broader applicability in pose-centric tasks. For instance, in blind inverse problems—certain parameters in $\mathcal{A}$ (e.g., camera models in HMR) are unknown—generative methods are less straightforward to implement. ZeDO [44], a recent study focusing on the 2D-3D lifting task, adopts the ScoreSDE [8] framework and refines camera translations by solving an optimization sub-problem after each generative step. However, directly porting this strategy to HMR is non-trivial, owing to the added complexity of body shape parameter optimization—a feature currently absent in our DPoser model. Although some state-of-the-art techniques [25, 24] offer solutions by jointly modeling operator $\mathcal{A}$ and data distributions, a full-fledged discussion on this subject is beyond this paper’s purview and remains an open question for future work.
F More Qualitative Results
We show more qualitative results for pose generation (Fig. S-3), pose completion (Fig. S-4), human mesh recovery (Fig. S-5), and motion denoising (Fig. S-6).
Figure S-3: Pose generation. DPoser can generate diverse and realistic poses.
Figure S-4: Pose completion. (a) Right arm under occlusion. (b) Trunk under occlusion.
Figure S-5: Human mesh recovery. (a) Initialization using mean poses and default camera. *Ground truth for the EHF dataset is annotated in SMPL-X [35], which extends SMPL [33] with fully articulated hands and an expressive face. (b) Initialization using the CLIFF [51] prediction.
[…]

### 23 Mar 2024 v2

@@
 Abstract
-Modeling human pose is a cornerstone in applications from human-robot interaction to augmented reality, yet crafting a robust human pose prior remains a challenge due to biomechanical constraints and diverse human movements.
-Traditional priors like VAEs and NDFs often fall short in realism and generalization, especially in extreme conditions such as unseen noisy poses.
+This work targets to construct a robust human pose prior.
+However, it remains a persistent challenge due to biomechanical constraints and diverse human movements.
+Traditional priors like VAEs and NDFs often exhibit shortcomings in realism and generalization, notably with unseen noisy poses.
 To address these issues, we introduce DPoser, a robust and versatile human pose prior built upon diffusion models.
-Designed with optimization frameworks, DPoser seamlessly integrates into various pose-centric applications, including human mesh recovery, pose completion, and motion denoising.
-Specifically, by formulating these tasks as inverse problems, we employ variational diffusion sampling for efficient solving.
-Furthermore, acknowledging the disparity between the articulated poses we focus on and structured images in previous research, we propose a truncated timestep scheduling to boost performance on downstream tasks.
-Our exhaustive experiments demonstrate DPoser’s superiority over existing state-of-the-art pose priors across multiple tasks.
+DPoser regards various pose-centric tasks as inverse problems and employs variational diffusion sampling for efficient solving.
+Accordingly, designed with optimization frameworks, DPoser seamlessly benefits human mesh recovery, pose generation, pose completion, and motion denoising tasks.
+Furthermore, due to the disparity between the articulated poses and structured images, we propose truncated timestep scheduling to enhance the effectiveness of DPoser.
+Our approach demonstrates considerable enhancements over common uniform scheduling used in image domains, boasting improvements of 5.4%, 17.2%, and 3.8% across human mesh recovery, pose completion, and motion denoising, respectively.
+Comprehensive experiments demonstrate the superiority of DPoser over existing state-of-the-art pose priors across multiple tasks.
 Corresponding authors: Yulun Zhang and Haoqian Wang.
+Keywords:
+Human Pose Prior, Diffusion Model
+1 Introduction
+Accurate modeling of human pose is a fundamental research topic that can benefit various applications, from human-robot interaction to augmented and virtual reality experiences.
+Many real-world applications rely on a prior distribution of valid human poses to perform tasks like body model fitting, motion capture, and gesture recognition.
+The complexity of human biomechanics, coupled with the extensive kinematic variability in movement patterns, presents a significant challenge in constructing a robust and realistic human pose prior.
+Previous efforts to model human pose prior have mainly employed techniques such as Gaussian Mixture Models (GMMs) [1], Variational Autoencoders (VAEs) [37], and Neural Distance Fields (NDFs) [49].
+Each technique, however, faces its own set of limitations.
+GMMs, for instance, might lead to the generation of implausible poses due to their unbounded nature.
+VAEs, restricted by their Gaussian assumptions, tend to generate average poses that may not accurately capture the full spectrum of human actions.
+Meanwhile, NDFs have shown promise in 3D surface modeling but struggle with generalizing across the complex, high-dimensional landscape of human pose manifolds.
+These limitations highlight a pressing need for a more comprehensive and dependable approach to modeling human pose priors, an endeavor this work seeks to address.
+Recently, Diffusion models [16, 48, 11, 20] have gained traction for their prowess in capturing complex, high-dimensional data distributions and enabling versatile sampling techniques.
+Their application has been seen in generating lifelike human motion sequences [56, 42] and functioning as multi-hypothesis pose estimators from 2D inputs [17, 8].
+However, these models are designed for specific generation tasks or tailored to work with conditional input data, which limits their applicability in broader contexts.
+The potential of diffusion models as a universal human pose prior remains largely untapped, and effective optimization methods for diverse tasks remain unanswered.
 Figure 1: An overview of DPoser’s versatility and performance across multiple pose-related tasks.
 Built on diffusion models, DPoser serves as a robust and adaptable pose prior.
-Shown are applications in (a) pose generation, (b) human mesh recovery, (c) motion denoising, and (d) pose completion.
+Shown are scenarios in (a) pose generation, (b) human mesh recovery, (c) motion denoising, and (d) pose completion.
 DPoser consistently outstrips existing priors like VPoser [37] in performance benchmarks.
-1 Introduction
-The accurate modeling of human pose is an essential building block in a variety of applications, ranging from human-robot interaction to augmented and virtual reality experiences.
-Many real-world applications rely on a prior distribution of valid human poses to perform tasks like body model fitting, motion capture, and gesture recognition.
-Existing methods use Gaussian Mixture Models (GMMs) [34], Variational Autoencoders (VAEs) [35], or Neural Distance Fields (NDFs) [36] to model the complex human pose prior.
-However, these models have limitations.
-GMMs can generate implausible poses due to their unbounded nature, VAEs are constrained by Gaussian assumptions, and NDFs exhibit limited generalization capabilities, requiring additional training strategies for extreme poses.
-These shortcomings underscore the need for a more robust and reliable human pose prior, a gap that this work aims to fill.
-Diffusion models [10, 8, 12, 14] have recently gained traction for their prowess in capturing complex, high-dimensional data distributions and enabling versatile sampling techniques.
-Although they have been employed in specialized applications like generating realistic human motion sequences [45, 46] or serving as multi-hypothesis pose estimators based on 2D inputs [41, 43], their utility has generally been constrained.
-These models are designed for generation or tailored to work with specific types of input data, which limits their applicability in broader contexts.
-Thus far, the potential of diffusion models as a universal human pose prior remains largely untapped, and effective optimization methods for diverse applications remain unanswered.
 In this work, we propose DPoser, a novel approach that leverages time-dependent denoiser learned from expansive motion capture datasets to construct a robust human pose prior.
-Specifically, by viewing various pose-centric tasks as inverse problems, we propose to employ variational diffusion sampling techniques [18] to incorporate DPoser as a regularization term within the existing optimization framework e.g., SMPLify [34].
-Furthermore, our experimental investigations into the generative aspects of articulated human poses reveal that the major information during diffusion is concentrated on the “clean” end of the perturbed data, as opposed to the entire noise-to-clean trajectory.
-Motivated by this insight, we introduce a truncated annealed timestep scheduling for optimization, which demonstrates superior efficiency over conventional random and full annealed schedules across multiple downstream tasks.
-In summary, our main contributions of this work can be summarized as follows:
-  • We introduce DPoser, a novel framework based on diffusion models to craft a robust and flexible human pose prior, geared for seamless integration across diverse pose-related applications via test-time optimization.
-  • Utilizing a fresh perspective on inverse problems, we employ variational diffusion sampling during test-time optimization and introduce truncated timestep scheduling, showing superior efficiency over traditional methods.
-  • Through extensive experiments and ablation studies, we establish that DPoser not only outshines state-of-the-art (SOTA) pose priors in a variety of applications but also provides valuable insights into training techniques such as suitable rotation representations for peak performance.
-2 Related works
+We regard various pose-centric tasks as inverse problems and suggest the integration of DPoser via variational diffusion sampling techniques [33] as a regularization component within optimization frameworks like SMPLify [1].
+Furthermore, our investigations reveal that significant pose-related information during diffusion is predominantly located at the latter stages of the diffusion trajectory.
+This revelation inspired us to develop a novel truncated timestep scheduling strategy for optimization.
+Our method outperforms the standard uniform scheduling, showing gains of 5.4%, 17.2%, and 3.8% in human mesh recovery, pose completion, and motion denoising, respectively.
+In summary, our main contributions are as follows:
+  • We introduce DPoser, a novel framework based on diffusion models to craft a robust and flexible human pose prior, geared for seamless integration across diverse pose-related tasks via test-time optimization.
+  • We analyze the impact of diffusion timesteps in the pose domain and propose truncated scheduling for more efficient optimization.
+  • Through extensive experiments, we establish that DPoser outshines state-of-the-art (SOTA) pose priors in a variety of downstream tasks.
+2 Related Work
 2.1 Human Pose Priors
-Human body models such as SMPL [33] serve as powerful tools for parameterizing both pose and shape, thereby offering a comprehensive framework for describing a wide array of human gestures.
+Human body models such as SMPL [31] serve as powerful tools for parameterizing both pose and shape, thereby offering a comprehensive framework for describing human gestures.
 Within the SMPL model, body poses are captured using rotation matrices or joint angles linked to a kinematic skeleton.
 Adjusting these parameters enables the representation of a diverse range of human actions.
-However, due to inherent biomechanical constraints, plausible poses inherently reside in a lower-dimensional manifold.
-Numerous approaches [34, 35, 36] have been proposed to model this intrinsic pose prior.
-Classical generative models like GMMs, VAEs [16], and GANs [17] have been employed to capture the complex underlying distribution, demonstrating significant improvements in downstream tasks such as human mesh recovery [49, 50].
-Additionally, some research has explored task-specific conditional pose priors, augmenting the model with additional inputs like image features [42, 48], 2D joint coordinates [43], or preceding poses in motion sequences [40, 39].
-Our work adopts the unconditional pose prior framework.
-We train DPoser on large-scale motion capture datasets without auxiliary inputs like paired images or captions, to ensure its applicability across a broad spectrum of pose-related tasks.
+Nonetheless, feeding unrealistic poses into these models can result in non-viable human figures, primarily because plausible human poses are confined within a complex, high-dimensional manifold due to biomechanical constraints.
+Various strategies [1, 37, 49, 9] have been put forward to build human pose priors.
+Generative frameworks like GMMs, VAEs [22], and Generative Adversarial Networks (GANs) [13] have shown promise in encapsulating the multifaceted pose distribution, facilitating advancements in tasks like human mesh recovery [19, 12].
+Further, some studies have delved into conditional pose priors tailored to specific tasks, incorporating extra information such as image features [39, 3], 2D joint coordinates [8], or sequences of preceding poses [28, 40].
+Our initiative leans towards an unconditional pose prior approach, training DPoser on extensive motion capture data without relying on additional inputs like images or text, aiming for a versatile application across various pose-related scenarios.
 2.2 Diffusion Models for Pose-centric Tasks
@@
 In a similar vein, DiffusionPose [39] and GFPose [8] employ the generation-based pipeline but take different approaches in conditioning.
-ZeDO [44] focuses on 2D-to-3D pose lifting, while works like Diff-HMR [48] and DiffHand [47] extend to estimating SMPL parameters and hand mesh vertices.
-BUDDI [62] stands out for using diffusion models to capture the joint distribution of interacting individuals and leveraging SDS loss [27, 28] for test-time optimization.
-Our work, DPoser, parallels BUDDI in similar optimization implementation.
-However, we distinguish ourselves by introducing a wider perspective of inverse problems and proposing advanced timestep scheduling based on characteristics of human poses.
-Moreover, DPoser is designed to work as a flexible regularization term, making it adaptable to diverse pose-centric tasks.
-It should also be noted that unlike most works [44, 42, 43, 41] that focus on joint positions, DPoser targets the more complex issue of joint rotations, adding an extra layer of challenge due to the intricacy of rotation representations.
-3 Preliminary: Score-based Diffusion Models
+Further, ZeDO [18] concentrates on 2D-to-3D pose lifting, while Diff-HMR [3] and DiffHand [24] explore estimating SMPL parameters and hand mesh vertices, respectively.
+BUDDI [34] stands out for using diffusion models to capture the joint distribution of interacting individuals and leveraging SDS loss [38, 52] for optimization during testing phases.
+While DPoser shares a similar optimization implementation with BUDDI, it sets itself apart by introducing a wider perspective of inverse problems and equipping an innovative timestep scheduling strategy tailored to the characteristics of human poses.
+Unlike other approaches [18, 39, 8, 17] that primarily focus on 3D location-based representation, DPoser takes on the more demanding task of modeling SMPL-based rotation pose representation.
+This adds complexity due to the intricacies involved in representing rotations, positioning DPoser as a more versatile solution within the realm of pose-centric tasks.
+3 Methods
+3.1 Preliminary: Score-based Diffusion Models
 Diffusion models [43, 47, 48, 16] operationalize generative processes by inverting a predefined forward diffusion process, typically formulated as a linear stochastic differential equation (SDE).
 Formally, the data trajectory $\left\{\mathbf{x}(t)\in\mathbb{R}^{n}\right\}_{t\in[0,1]}$ follows the forward SDE given by:
-$\mathrm{d}\mathbf{x}=\mu(t)\mathbf{x}\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}.$ (1)
-Here, $\mu(t)\mathbf{x}\in\mathbb{R}^{n}$ and $g(t)\in\mathbb{R}$ represent the drift and diffusion coefficients, while $\mathbf{w}$ is a standard Wiener process.
+$\mathrm{d}\mathbf{x}=\mu(t)\mathbf{x}\mathrm{d}t+g(t)\mathrm{d}\mathbf{w},$ (1)
+where $\mu(t)\mathbf{x}\in\mathbb{R}^{n}$ and $g(t)\in\mathbb{R}$ represent the drift and diffusion coefficients, while $\mathbf{w}$ is a standard Wiener process.
 The affine drift coefficients ensure analytically tractable Gaussian perturbation kernels, denoted by $p_{0t}(\mathbf{x}_{t}\mid\mathbf{x})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x},\sigma_{t}^{2}\mathbf{I})$, where the exact coefficients $\alpha_{t},\sigma_{t}$ can be obtained with standard techniques [41].
@@
 The so-called score function [29], $\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\right)$, serves as an unknown term in Eq.
-(2) and can be approximated by a neural network parameterized as $\epsilon_{\theta}(\mathbf{x}_{t};t)\approx-\sigma_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\right)$ [^1] This parameterization is obtained from the deep connection between the noise prediction in diffusion models and score function estimation in score-based models.
+(2) and can be approximated by a neural network parameterized as $\epsilon_{\phi}(\mathbf{x}_{t};t)\approx-\sigma_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\right)$ [^1] This parameterization is obtained from the deep connection between the noise prediction in diffusion models and score function estimation in score-based models.
 We provide a brief recap in the Appendix..
-To learn the score functions, employing denoising score matching techniques [2], we first introduce noise to the data points as per:
+To learn the score functions, employing denoising score matching techniques [50], we perturb the data points with noise as per:
 $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\epsilon,\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (3)
-Subsequently, we train the time-dependent noise predictor $\epsilon_{\theta}(\mathbf{x}_{t};t)$ using an L2-loss defined as [10]:
-$\mathbb{E}_{\mathbf{x}_{0}\sim p_{\mathrm{data}},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[w(t)||\epsilon-\epsilon_{\theta}(\mathbf{x}_{t};t)||_{2}^{2}\right],$ (4)
+Subsequently, feeding $\mathbf{x}_{t}$ and $t$ as input, we train the time-dependent noise predictor $\epsilon_{\phi}(\mathbf{x}_{t};t)$ using an L2-loss defined as [16]:
+$\mathbb{E}_{\mathbf{x}_{0}\sim p_{\mathrm{data}},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t\sim\mathcal{U}[0,1]}\left[w(t)||\epsilon-\epsilon_{\phi}(\mathbf{x}_{t};t)||_{2}^{2}\right],$ (4)
 where $w(t)$ denotes a positive weighting function.
@@
 Through techniques like Euler-Maruyama discretization, we can generate novel samples by simulating the reverse SDE.
-4 Methods
-4.1 Learning Pose Prior with Diffusion Models
-We propose to model pose prior based on the SMPL body model [33], which can be viewed as a differentiable function $[J,V]=M(\theta,\beta)$ that maps body joint angles $\theta\in\mathbb{R}^{3\times 21}$ and shape parameters $\beta\in\mathbb{R}^{10}$ to mesh vertices $V\in\mathbb{R}^{3\times 6890}$ and joint positions $J\in\mathbb{R}^{3\times 22}$.
+3.2 Learning Pose Prior with Unconditional Diffusion Models
+SMPL-based pose representation.
+To build a flexible 3D human pose prior, we propose to utilize the SMPL body model [31], which can be viewed as a differentiable function $[J,V]=M(\theta,\beta)$ that maps body joint angles $\theta\in\mathbb{R}^{3\times 21}$ and shape parameters $\beta\in\mathbb{R}^{10}$ to mesh vertices $V\in\mathbb{R}^{3\times 6890}$ and joint positions $J\in\mathbb{R}^{3\times 22}$.
 Our target is to model the distribution of joint angles $p(\theta)$.
-To achieve this, we train DPoser on mocap datasets, employing sub-VP SDEs introduced in ScoreSDE [8], defined as:
-$\mathrm{d}\mathbf{x}=-\frac{1}{2}\beta(t)\mathbf{x}\mathrm{d}t+\sqrt{\beta(t)(1-e^{-2\int_{0}^{t}\beta(s)\mathrm{d}s})}\mathrm{d}\mathbf{w},$ (5)
-where $\beta(t)$ denotes linear scheduled noise scales.
+Training of unconditional diffusion models.
+To this end, we adopt an unconditional diffusion model to learn the pose representation $\theta$.
+This approach aligns with a task-agnostic strategy, focusing solely on the distribution of 3D poses.
+We employ sub-VP SDEs as outlined in [48], which have demonstrated efficacy in sampling quality, for constructing our diffusion model.
+Specifically, our chosen forward SDE (Eq.
+(1)) is given by:
+$\mathrm{d}\mathbf{x}=-\frac{1}{2}\xi(t)\mathbf{x}\mathrm{d}t+\sqrt{\xi(t)(1-e^{-2\int_{0}^{t}\xi(s)\mathrm{d}s})}\mathrm{d}\mathbf{w},$ (5)
+where $\xi(t)$ denotes linear scheduled noise scales.
 The coefficients needed in Eq.
-(3) can be obtained as $\alpha_{t}=e^{-\frac{1}{2}\int_{0}^{t}\beta(s)\mathrm{d}s},\sigma_{t}=[1-e^{-\int_{0}^{t}\beta(s)\mathrm{d}s}]^{2}$.
-With the constructed forward process (Eq.
-(3)) to sample $\mathbf{x}_{t}$, we then apply the objective in Eq.
-(4) to train the noise predictor $\epsilon_{\theta}(\mathbf{x}_{t};t)$ with weights $w(t)=\sigma_{t}^{2}$ as suggested in ScoreSDE [8].
-In our experiments, we find the selection of rotation representations, such as axis-angle and 6D rotations [59], to be crucial for optimal performance when coupled with suitable normalization.
-Despite experimenting with incorporating body models via auxiliary losses, as in prior work [35, 39], we note that it imposes a computational burden without evident gains in downstream task performance.
-Further methodological nuances are detailed in the Appendix.
+(3) can be obtained as $\alpha_{t}=e^{-\frac{1}{2}\int_{0}^{t}\xi(s)\mathrm{d}s},\sigma_{t}=1-e^{-\int_{0}^{t}\xi(s)\mathrm{d}s}$.
+During training, we initiate with a clean data point $\mathbf{x}_{0}$—essentially, our pose representation $\theta$—and introduce noise to generate samples $\mathbf{x}_{t}$ according to the forward process detailed in Eq.
+(3).
+Then we apply the objective in Eq.
+(4) to train the noise predictor $\epsilon_{\phi}(\mathbf{x}_{t};t)$ with weights $w(t)=\sigma_{t}^{2}$ as suggested in [48].
 3.3 Optimization Leveraging Diffusion Priors
-Figure 2: Overview of the DPoser Methodology.
-Panel (a) presents three tasks: human mesh recovery, pose completion, and motion denoising, with omissions like camera optimization and bending loss for clarity.
-Panel (b) demonstrates the DPoser regularization process, introducing noise to the current pose and applying a one-step denoiser to achieve a denoised pose.
-$L_{\text{DPoser}}$ is computed between the denoised and current pose.
-Panel (c) outlines the optimization process from initial to current and fitted poses via loss minimization.
-The acquired score functions or noise predictors, denoted as $\epsilon_{\theta}(\mathbf{x}_{t};t)$, permit the direct generation of plausible poses through Eq.
+The acquired score functions or noise predictors, denoted as $\epsilon_{\phi}(\mathbf{x}_{t};t)$, permit the direct generation of plausible poses through Eq.
 (2).
@@
 Inverse problem formulation. Consider an original signal $\mathbf{x}_{0}$. Inverse problems can be encapsulated by Eq. (6) as:
-$\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})+\mathbf{n},\quad\mathbf{y},\mathbf{n}\in\mathbb{R}^{d},~{}\mathbf{x}_{0}\in\mathbb{R}^{n},$ (6)
-where $\mathcal{A}$ symbolizes the measurement operator and $\mathbf{n}$ constitutes noise, assumed to be white Gaussian $\mathcal{N}(\mathbf{0},\sigma_{n}^{2}\mathbf{I})$ in our context.
-As an illustration, inverse kinematics can be considered as inferring joint rotations (or body poses in SMPL [33]) from observed joint positions.
-Moreover, the 2D-3D lifting task can be modeled as a linear inverse problem, with $\mathcal{A}$ representing a camera projection model.
-To recover the original signal $\mathbf{x}_{0}$, in the Bayesian framework, our target is to sample from the posterior distribution $p\left(\mathbf{x}_{0}\mid\mathbf{y}\right)$.
-Diffusion-based inverse problem solving.
-Various techniques [26, 21, 19, 20, 22, 18] have been explored to simulate this posterior sampling process based on diffusion priors.
-For some noiseless linear inverse problems, such as completion, alternating projections onto measurement subspaces are commonly employed [8, 21, 20].
-For instance, in hand pose completion, a standard iterative generation process using diffusion models can be applied, replacing known parts of the pose at each iteration.
-However, such projections are inconvenient to build general optimization frameworks concerned in our work.
-To navigate these challenges, we adopt variational diffusion sampling [18], efficient in computation and straightforward in implementation.
+$\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})+\mathbf{n},\quad\mathbf{y},\mathbf{n}\in\mathbb{R}^{d},~{}\mathbf{x}_{0}\in\mathbb{R}^{n},\vspace{-2mm}$ (6)
+where $\mathcal{A}$ symbolizes the measurement operator and $\mathbf{n}$ constitutes noise, assumed to be white Gaussian $\mathcal{N}(\mathbf{0},\sigma_{n}^{2}\mathbf{I})$.
+In the context targeted in this study, $\mathbf{x}_{0}$ always refers to body poses in SMPL [31].
+This formulation allows us to approach various pose-centric tasks by adapting $\mathcal{A}$ and interpreting $\mathbf{y}$ accordingly:
+  • Pose completion: Here, $\mathcal{A}$ serves as a mask matrix to simulate partially observed poses, with $\mathbf{y}$ being the incomplete pose data.
+  • Motion denoising: In this scenario, $\mathcal{A}$ applies SMPL’s forward kinematics, treating $\mathbf{y}$ as the observed noisy 3D joints.
+  • Human mesh recovery: $\mathcal{A}$ integrates SMPL’s forward kinematics and camera projection to relate $\mathbf{y}$ to 2D joint observations in images.
+The aim is to recover the original signal $\mathbf{x}_{0}$, where, within the Bayesian framework, our objective shifts to sampling from the posterior distribution $p\left(\mathbf{x}_{0}\mid\mathbf{y}\right)$.
+Solving inverse problems with diffusion models.
+Various techniques [14, 21, 6, 7, 45, 33] have been explored to simulate this posterior sampling process based on unconditional diffusion priors $p\left(\mathbf{x}_{0};\phi\right)$.
+Among them, the sampling-based scheme is widely explored and applied in tasks like image restoration.
+These methods incorporate the observation information $\mathbf{y}$ into the generation process of $\mathbf{x}_{0}$ through techniques like gradient guidance [6, 7] and back projection [48, 21, 7].
+However, such methods rooted in generation are inconvenient for handling diverse pose-related tasks.
+To navigate these challenges, we adopt variational diffusion sampling [33] to build general optimization frameworks.
 Specifically, it employs a variational distribution $q\left(\mathbf{x}_{0}\mid\mathbf{y}\right):=\mathcal{N}(\mu,\sigma^{2}\mathbf{I})$ and aims to minimize the Kullback-Leibler (KL) divergence between this variational distribution and the true posterior, mathematically expressed as $KL\big{(}q\left(\mathbf{x}_{0}\mid\mathbf{y}\right)\parallel p\left(\mathbf{x}_{0}\mid\mathbf{y}\right)\big{)}$.
-Informed by existing theorems concerning the likelihood properties of diffusion models [9], and under the assumption of zero variance ($\sigma\approx 0$), the optimization problem of seeking $\mathbf{x}_{0}$ (i.e., $\mu$) can be formulated as minimizing [18]:
-$\|\mathbf{y}-\mathcal{A}(\mathbf{x}_{0})\|^{2}+w_{t}(\mathtt{sg}[\epsilon_{\theta}(\mathbf{x}_{t};t)-\epsilon])^{\top}\mathbf{x}_{0},$ (7)
+Further, under the assumption of zero variance ($\sigma\approx 0$), the optimization problem of seeking $\mathbf{x}_{0}$ (i.e., $\mu$) can be formulated as minimizing [46, 33]:
+$\|\mathbf{y}-\mathcal{A}(\mathbf{x}_{0})\|^{2}+w_{t}(\mathtt{sg}[\epsilon_{\phi}(\mathbf{x}_{t};t)-\epsilon])^{\top}\mathbf{x}_{0},\vspace{-1.5mm}$ (7)
 where $w_{t}$ denotes the loss weights and $\epsilon$ is sampled from the standard Gaussian distribution.
@@
 (3), resulting in $\mathbf{x}_{t}$.
-Subsequently, the gradients $[\epsilon_{\theta}(\mathbf{x}_{t};t)-\epsilon]$ are applied to the optimization variable $\mathbf{x}_{0}$.
+Subsequently, the gradients $[\epsilon_{\phi}(\mathbf{x}_{t};t)-\epsilon]$ are applied to the optimization variable $\mathbf{x}_{0}$.
 In a nutshell, this framework [33] provides a flexible yet robust strategy for employing diffusion priors in generic optimization problems, serving as a cornerstone for our work.
-DPoser regularization.
-To shed more light on the working mechanism, we reformulate the regularization term as [^2] We detail another perspective of Score Distillation Sampling (SDS) [27, 28] to understand our DPoser regularization in the Appendix.:
+Figure 2: Overview of the DPoser Methodology.
+Panel (a) presents three tasks: human mesh recovery, pose completion, and motion denoising, with omissions like camera optimization for clarity.
+Panel (b) demonstrates the DPoser regularization process, introducing noise to the current pose and applying a one-step denoiser to achieve a denoised pose.
+$L_{\text{DPoser}}$ is computed between the denoised and current pose.
+Panel (c) outlines the optimization process from initial to fitted poses via loss minimization.
+Introducing DPoser regularization. To shed more light on the working mechanism, we propose to reformulate the regularization term as:
 $L_{\mathrm{DPoser}}$ $=w_{t}||\mathbf{x}_{0}-\mathtt{sg}[\mathbf{\hat{x}}_{0}(t)]||_{2}^{2},\text{where}$ (8)
-$\mathbf{\hat{x}}_{0}(t)$ $=\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\theta}(\mathbf{x}_{t};t)}{\alpha_{t}}.$ (9)
-Here, $\mathbf{\hat{x}}_{0}(t)$ serves as a one-step denoised estimate, leveraging the trained diffusion model $\epsilon_{\theta}(\mathbf{x}_{t};t)$.
-We show that the gradient of our DPoser regularization aligns with Eq.
-(7).
+$\mathbf{\hat{x}}_{0}(t)$ $=\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\phi}(\mathbf{x}_{t};t)}{\alpha_{t}}.$ (9)
+Here, $\mathbf{\hat{x}}_{0}(t)$ functions as a precise one-step denoising prediction using the diffusion model $\epsilon_{\phi}(\mathbf{x}_{t};t)$.
+This approach effectively encourages the current pose $\mathbf{x}_{0}$ towards a denoised, plausible pose distribution, employing a straightforward L2-loss within the DPoser regularization framework.
+Further, the theoretical foundation of our regularization demonstrates its alignment with the gradient direction of variational diffusion sampling (Eq.
+(7)).
 Proof: Differentiating Eq. (8) with respect to $\mathbf{x}_{0}$ yields:
 $\nabla_{\mathbf{x}_{0}}L_{\mathrm{DPoser}}$ $=2w_{t}(\mathbf{x}_{0}-\mathbf{\hat{x}}_{0}(t))$
-$=2w_{t}(\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon}{\alpha_{t}}-\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\theta}(\mathbf{x}_{t};t)}{\alpha_{t}})$
-$=2w_{t}\frac{\sigma_{t}}{\alpha_{t}}(\epsilon_{\theta}(\mathbf{x}_{t};t)-\epsilon)$
-$\propto(\epsilon_{\theta}(\mathbf{x}_{t};t)-\epsilon).$ (10)
-Therefore, our DPoser regularization $L_{\mathrm{DPoser}}$ can be implemented as an L2-loss.
-In this view, during optimization, DPoser works by pulling the current pose $\mathbf{x}_{0}$ closer to the learned plausible pose distribution through denoising.
+$=2w_{t}(\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon}{\alpha_{t}}-\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\phi}(\mathbf{x}_{t};t)}{\alpha_{t}})$
+$=2w_{t}\frac{\sigma_{t}}{\alpha_{t}}(\epsilon_{\phi}(\mathbf{x}_{t};t)-\epsilon)$
+$\propto(\epsilon_{\phi}(\mathbf{x}_{t};t)-\epsilon).$ (10)
+Thus, $L_{\mathrm{DPoser}}$ represents a more intuitive approach to variational diffusion sampling.
+By incorporating alongside task-specific loss functions, this regularization term enhances the plausibility of the resultant poses.
 DPoser across pose-related tasks.
-Capitalizing on DPoser’s inherent flexibility, we adapt it to various human pose tasks, with a spotlight on human mesh recovery as illustrated in Fig.
+DPoser excels in versatility, enabling its seamless application in a spectrum of human pose-related tasks.
+Its adaptability is especially evident in our human mesh recovery approach, as depicted in Fig.
 2.
-For an exhaustive examination of DPoser’s utility across different pose-related applications including pose completion and motion denoising, we direct the reader to our Appendix.
-By adapting the objective function from the SMPLify framework [34], we supplant the originally used GMM prior with our own DPoser regularization term, $L_{\mathrm{DPoser}}$, while also forgoing the intricate interpenetration error term.
-The optimization objective, incorporating both pose $\theta$ and shape $\beta$ parameters in SMPL [33], is articulated as:
+For an exhaustive examination of DPoser’s utility across tasks like pose completion and motion denoising, we direct the reader to our Appendix.
+Human mesh recovery aims to deduce the human pose and shape from single-image inputs.
+In this context, we refine the optimization function derived from the SMPLify framework [1], integrating DPoser as a regularization term, $L_{\mathrm{DPoser}}$, and streamlining the process by omitting the intricate interpenetration error component.
+The modified optimization objective, engaging both pose $\theta$ and shape $\beta$ parameters from the SMPL model [31], is defined as:
 $L(\theta,\beta)=L_{J}+w_{\theta}L_{\theta}+w_{\beta}L_{\beta}+w_{\alpha}L_{\mathrm{DPoser}}.$ (11)
-Herein, the reprojection loss $L_{J}$ serves as the data fidelity term, mathematically expressed as:
+The reprojection loss $L_{J}$, acting as the data fidelity measure, is defined by:
 $L_{J}$ $=\sum_{i\in\text{Joints}}\lambda_{i}\rho\left(\Pi_{C}\left(M_{J}(\theta,\beta)_{i}\right)-J^{\text{est}}_{i}\right),$ (12)
-where $M_{J}(\theta,\beta)$ and $\Pi_{C}$ symbolize the 3D joint positions regressed from SMPL and the 3D-to-2D camera projection function, respectively.
-$J^{\text{est}}$ are the 2D keypoints estimated using an off-the-shelf 2D pose estimator (OpenPose [60] in our experiments), and $\lambda_{i}$ indicates the confidence score of joint $i$.
-The Geman-McClure error function [61] $\rho$ serves to robustly quantify discrepancies in 2D joint positions.
-The bending term $L_{\theta}$ is incorporated to penalize excessive bending at the elbows and knees, formulated as $L_{\theta}=\sum_{i\in\text{(elbows, knees)}}\exp(\boldsymbol{\theta}_{i})$.
-The shape regularizer $L_{\beta}$ is a simple L2 norm, i.e., $L_{\beta}=\|\beta\|_{2}^{2}$.
-The constant weights for prior terms are denoted as $w_{\theta},w_{\beta}$ and $w_{\alpha}$, respectively.
-Contrasting SMPLify’s GMM-based prior, we employ $L_{\mathrm{DPoser}}$ to draw the pose closer to our learned manifold.
+where $M_{J}(\theta,\beta)$ calculates the 3D joint coordinates through SMPL’s forward kinematics.
+The function $\Pi_{C}$ maps these 3D coordinates into 2D space, aligning with the camera’s perspective.
+$J^{\text{est}}$ refers to the 2D keypoints estimated using an off-the-shelf 2D pose estimator (in our case, ViTPose [55]), with $\lambda_{i}$ reflecting the confidence score for each joint $i$.
+The Geman-McClure error function ($\rho$) is employed to assess the discrepancy in 2D joint locations reliably.
+To mitigate the issue of overfitting, which often leads to unrealistic poses when solely minimizing reprojection loss, several regularization terms are introduced.
+Specifically, alongside our body prior $L_{\mathrm{DPoser}}$, the bending term $L_{\theta}$ is incorporated to penalize excessive bending at the elbows and knees, formulated as $L_{\theta}=\sum_{i\in\text{(elbows, knees)}}\exp(\boldsymbol{\theta}_{i})$.
+Additionally, the shape regularization term $L_{\beta}=\|\beta\|_{2}^{2}$ is employed to maintain the body shape within plausible bounds.
+The weights for prior terms are denoted as $w_{\theta},w_{\beta}$ and $w_{\alpha}$, respectively.
 Given the structure of $L_{\mathrm{DPoser}}$ (as seen in Eq.
 (8)), a crucial aspect lies in judiciously selecting the diffusion timestep $t$ during the iterative optimization process.
-We address this concern by introducing our novel truncated timestep scheduling strategy in the subsequent section.
+In the subsequent section, we address this concern by introducing our novel truncated timestep scheduling strategy.
 3.4 Test-time Truncated Timestep Scheduling
+Motivation from pose generation.
+Adapting techniques from the image domain to pose data requires a nuanced understanding of the differences between the two.
+Previous image-based research [4] shows that initial timesteps (larger $t$) correspond to the perceptual content, while later timesteps refine details.
+Pose data, however, lacks this structured layering and spatial redundancy, indicating a need for a tailored timestep approach in the diffusion process.
 Figure 3: Illustration of the rationale behind our proposed truncated timestep scheduling.
 We employ the deterministic DDIM sampler [44] with limited steps and assess the quality of generated poses using the Self-Intersection percentage (SI).
+As depicted in Fig.
+3, we find that pose generation doesn’t benefit from the early timesteps as image generation does.
+The significant stages of pose refinement occur at smaller $t$, specifically when $t\leq 0.3$.
+A uniform distribution of timesteps, as tested in (b) with only five steps, proves less effective for pose data.
+In contrast, allocating these steps toward the latter end of the diffusion process, as in (c), yields significantly better samples, implying the critical information is not evenly distributed but rather is concentrated toward the end.
+Truncated timestep scheduling.
+Based on these insights, we propose a shift from standard uniform timestep sampling to a truncated strategy, especially for pose data.
+By focusing on the last timesteps, particularly between 0.2 and 0.0, we target the interval rich in pose-specific information.
+Specifically, based on the linear descending scheduling, the truncated timestep $t$ for each optimization step can be expressed as:
+$t=t_{\text{max}}-\frac{(t_{\text{max}}-t_{\text{min}})\times\text{iter}}{N-1}.$ (13)
+where $N$ denotes the total number of optimization iterations, and iter signifies the current iteration.
+This formulation is integral to our proposed optimization framework, which is comprehensively summarized in Algorithm 1.
+The practical implementation typically involves setting the truncated range to $[0.2,0.05]$.
 Algorithm 1 Test-time Optimization with DPoser
-1:A trained diffusion model $\epsilon_{\theta}(\mathbf{x}_{t};t)$, task-specific loss $L_{\text{task}}$, range of diffusion timesteps $[t_{\text{max}},t_{\text{min}}]$, number of optimization iterations $N$, DPoser weights $w_{t}$
-2:Initialization of parameters $\mathbf{x}_{0}$
+1:A trained diffusion model $\epsilon_{\phi}(\mathbf{x}_{t};t)$, task-specific loss $L_{\text{task}}$, range of diffusion timesteps $[t_{\text{max}},t_{\text{min}}]$, number of optimization iterations $N$.
+2:Initialization of SMPL body pose parameters $\mathbf{x}_{0}$
 3:for $\text{iter}=0,1,\ldots,N-1$ do
@@
 6:     $\mathbf{x}_{t}\leftarrow\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\epsilon$triangleright Forward diffusion
-7:     $\mathbf{\hat{x}}_{0}(t)\leftarrow\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\theta}(\mathbf{x}_{t};t)}{\alpha_{t}}$triangleright One-step denoiser
-8:     $L_{\text{DPoser}}\leftarrow w_{t}\lVert\mathbf{x}_{0}-\text{sg}[\mathbf{\hat{x}}_{0}(t)]\rVert_{2}^{2}$
+7:     $\mathbf{\hat{x}}_{0}(t)\leftarrow\frac{\mathbf{x}_{t}-\sigma_{t}\epsilon_{\phi}(\mathbf{x}_{t};t)}{\alpha_{t}}$triangleright One-step denoiser
+8:     $L_{\text{DPoser}}\leftarrow w_{t}\lVert\mathbf{x}_{0}-\text{sg}[\mathbf{\hat{x}}_{0}(t)]\rVert_{2}^{2}$triangleright DPoser regularization
 9:     $L_{\text{total}}\leftarrow L_{\text{task}}+L_{\text{DPoser}}$
@@
 12:return $\mathbf{x}_{0}$
-Insights from pose generation.
-In optimizing pose-related tasks, one fundamental challenge arises from the gap between the image-based data and the articulated pose parameters we manipulate.
-Prior studies in image domains [15] have noted that earlier timesteps (i.e., smaller $t$) are typically aligned with perceptual content, while the later timesteps are more closely associated with the refinement of details in the diffusion process.
-As illustrated in Fig.
-3, our experimental foray into employing diffusion models for pose generation yields an interesting insight into pose domains: the concentration of meaningful information resides at the “clean” side of the diffusion trajectory, rather than being dispersed uniformly from noise to clean states.
-This observation serves as the catalyst for our introduction of a truncated annealed timestep scheduling mechanism.
-Truncated timestep scheduling.
-Departing from traditional diffusion models that uniformly sample across the continuum from $1.0$ to $0.0$, we introduce a selective sampling strategy.
-Specifically, we truncate the timestep range to intervals such as $0.2\to 0$, thereby focusing on the data segments most laden with meaningful information.
-Based on the linear annealed scheduling, the truncated timestep $t$ for each optimization step can be formulated as:
-$t=t_{\text{max}}-\frac{(t_{\text{max}}-t_{\text{min}})\times\text{iter}}{N-1}.$ (13)
-Here, $N$ signifies the total optimization iterations, while iter represents the current optimization iteration.
-We summarize our DPoser optimization framework as Algorithm 1, where the truncated range $[t_{\text{max}},t_{\text{min}}]$ is typically set to $[0.2,0.001]$.
-Through ablation studies, we prove that the proposed truncated timestep scheduling offers significant advantages over the random, fixed, and full annealed ones utilized in previous works [62, 18, 19, 8], delivering more efficient optimization across a wide range of downstream tasks.
-This is achieved by concentrating computation on the most informative portions of the data trajectory, as informed by our preliminary sampling investigations on pose generation.
 4 Experiments
-In this section, we first elaborate on the implementation details and evaluation metrics.
-We showcase the robustness and versatility of DPoser across a spectrum of pose-centric applications, including pose generation, human mesh recovery, pose completion, and motion denoising.
-Additionally, we conduct an ablation study to delve into the key elements contributing to our model’s performance.
-Sample source APD $\uparrow$ SI $\downarrow$
-Real-world (AMASS) [52] 15.44 0.79
-GMM [34] 16.28 1.54
-VPoser [35] 10.75 1.51
-Pose-NDF [36] 18.75 1.97
-DPoser (ours) 14.28 1.21
-DPoser (ours)* 19.03 1.13
-Table 1: Comparative analysis of pose generation metrics.
-The discrepancy between visual impressions and APD/SI metrics is discussed, with reference to Fig.
-4.
-*Indicates the use of a DDIM sampler [11] with a reduced 10-step discretization.
-(a) AMASS samples
-(b) DPoser (ours)
-(c) DPoser (ours)*
-(d) GMM
-(e) Pose-NDF
-(f) VPoser
-Figure 4: Qualitative comparison of generated human poses: (b) illustrates naturalistic poses closely aligned with real-world data, whereas (c) shows poses that, despite superior APD and SI metrics, lack natural appearance.
-*We use a DDIM sampler [11] with only 10 steps.
+In this section, we showcase the robustness and versatility of DPoser across a spectrum of pose-centric tasks, including pose generation, human mesh recovery, pose completion, and motion denoising.
+Due to the page limit, we leave experimental details and more qualitative assessments in the Appendix.
 4.1 Experimental Setup
@@
 It draws inspiration from GFPose [8] but omits conditional input pathways for our unconditional setting.
-To stabilize training, we use an exponential moving average with a decay factor of 0.9999, as advised by ScoreSDE [8].
+To stabilize training, we use an exponential moving average with a decay factor of 0.9999, as advised by [48].
 The Adam optimizer, a learning rate of $2\times 10^{-4}$, and a batch size of 1280 govern the optimization process.
 The training of 800,000 iterations takes roughly 8 hours on a single Nvidia RTX 3090Ti GPU.
-Evaluation metrics. We adopt task-specific metrics:
-  • Pose Generation: Diversity and fidelity are evaluated using Average Pairwise Distance (APD) [37] and Self-Intersection rates (SI) [36], respectively.
-  • Human Mesh Recovery: Metrics include Mean Per Joint Position Error (MPJPE) and Procrustes-aligned MPJPE (PA-MPJPE).
-  • Pose Completion: We utilize both MPJPE and Mean Per-Vertex Position Error (MPVPE), for masked areas.
-  • Motion Denoising: Metrics are MPVPE and MPJPE.
+Evaluation metrics. To comprehensively evaluate our models across various tasks, following Pose-NDF [49], we adopt task-specific metrics:
+  • Pose Generation: Diversity and fidelity are evaluated using Average Pairwise Distance (APD) and Self-Intersection rates (SI), respectively.
+  • Human Mesh Recovery: The Procrustes-aligned Mean Per Joint Position Error (PA-MPJPE) measures the accuracy of recovered human meshes.
+  • Pose Completion: The Mean Per Joint Position Error (MPJPE) for masked body joints serves as the metric, focusing on the inferred occluded parts.
+  • Motion Denoising: Both MPJPE and the Mean Per-Vertex Position Error (MPVPE) are calculated to assess the denoising effectiveness.
 All errors are reported in millimeter units.
 4.2 Pose Generation
+(a) GAN-S [9]
+(b) DPoser (ours)
+(c) DPoser (ours)*
+(d) GMM [1]
+(e) Pose-NDF [49]
+(f) VPoser [37]
+Figure 4: Qualitative comparison of generated human poses: (b) illustrates naturalistic poses aligned with real-world data, whereas (c) shows poses that, despite superior metrics, lack natural appearance.
+*We use a DDIM sampler [44] with only 10 steps.
+Sample source APD $\uparrow$ SI $\downarrow$
+Real-world (AMASS) [32] 15.44 0.79
+GMM [1] 16.28 1.54
+VPoser [37] 10.75 1.51
+Pose-NDF [49] 18.75 1.97
+GAN-S [9] 15.68 1.27
+DPoser (ours) 14.28 1.21
+DPoser (ours)* 19.03 1.13
+Table 1: Comparative analysis of pose generation metrics.
+The discrepancy between visual impressions and APD/SI metrics is discussed, with reference to Fig.
+4.
+*Indicates the use of a reduced 10-step sampler.
 To commence, we delve into the capabilities of our DPoser model by generating samples from the learned manifold.
@@
 In summary, our findings underscore the need for a balanced evaluation strategy that merges quantitative metrics with qualitative observations.
-Beyond generation, DPoser also performs well in interpolation tasks, based on corresponding ODEs with marginal densities akin to Eq.
-(2[8].
-Additional experiments are detailed in the Appendix.
-Methods PA-MPJPE $\downarrow$ MPJPE $\downarrow$
-GMM [34] 57.53 79.66
-VPoser [35] 61.10 85.25
-Pose-NDF [36] 61.67 85.09
-DPoser (ours) 57.28 84.79
-CLIFF [51] 57.30 92.86
-+ GMM [34] 55.54 78.32
-+ VPoser [35] 56.12 79.68
-+ Pose-NDF [36] 55.50 78.39
-+ DPoser (ours) 54.87 77.62
-Table 2: Performance comparison of human mesh recovery on the EHF dataset [35] using two initialization methods.
-Methods Occ. left leg Occ. legs Occ. arms Occ. trunk
-MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
-Pose-NDF [36] ($S=1$) 499.35 467.26 503.66 470.82 378.61 312.16 319.65 208.35
-Pose-NDF [36] ($S=1$${}^{\dagger}$ 597.72 558.45 606.68 566.01 370.3 283.43 346.18 228.54
-VPoser [35] ($S=1$) 465.98 437.92 464.38 436.06 387.14 293.17 298.63 195.16
-VPoser [35] ($S=1$${}^{\dagger}$ 194.44 177.01 199.78 183.34 205.17 160.60 75.42 48.30
-VPoser [35] ($S=5$) 458.92 431.31 456.01 427.90 374.32 283.90 288.96 188.26
-VPoser [35] ($S=10$) 455.44 427.79 450.88 423.04 371.26 281.62 286.55 186.28
-DPoser (ours) ($S=1$) 85.46 80.66 107.07 100.97 108.27 84.37 46.37 29.99
-DPoser (ours) ($S=5$) 48.61 44.79 74.67 69.64 75.92 59.90 26.59 16.68
-DPoser (ours) ($S=10$) 39.77 36.15 65.24 60.54 66.77 52.75 21.58 13.41
-Table 3: Performance metrics for pose completion under varying occlusion scenarios.
-$S$ denotes the number of hypotheses.
-${}^{\dagger}$ Comparison methods are implemented in the optimization paradigm.
 4.3 Human Mesh Recovery
+Initialization No fitting GMM [1] VPoser [37] Pose-NDF [49] GAN-S [9] DPoser(Ours)
+from scratch 108.57 58.32 58.08 57.87 57.26 56.05
+CLIFF [25] 56.62 51.02 49.39 49.50 49.58 49.05
+Table 2: Performance comparison of human mesh recovery on the EHF dataset [37] using two initialization methods. PA-MPJPE is reported as the metric.
+Figure 5: Human mesh recovery.
+(a) Fitting from scratch.
+*Ground truth for the EHF dataset is annotated in SMPL-X [37], which extends SMPL [31] with fully articulated hands and an expressive face.
+(b) Initialization using the CLIFF [25] prediction.
 (a) EHF dataset
@@
 (d) UBody dataset
-Figure 5: Qualitative evaluations of HMR leveraging DPoser as pose prior. CLIFF [51] serves as the optimization initializer.
-We probe the efficacy of DPoser in HMR, focusing on estimating human pose and shape from monocular images.
+Figure 6: Qualitative evaluations of human mesh recovery leveraging DPoser as pose prior. CLIFF [25] serves as the optimization initializer.
+We probe the efficacy of DPoser in human mesh recovery (HMR), focusing on estimating human pose and shape from monocular images.
 We conduct experiments on the EHF dataset [37] and benchmark our method against existing SOTA priors.
 Our optimization-based framework incorporates two initialization paradigms: (1) a baseline initialization that utilizes mean pose values and a default camera setup, and (2) an advanced initialization scheme that leverages CLIFF [25], a pre-trained regression-based model tailored for HMR.
-As shown in Tab.
-2, DPoser demonstrates robust performance in HMR.
-With baseline initialization, it closely rivals the strong constraints of the GMM prior [34].
-When enhanced by advanced initialization via CLIFF, DPoser surpasses established deep priors like Pose-NDF [36].
-Fig.
-5 visually confirms DPoser’s superior efficacy and adaptability across multiple datasets including EHF [35], MSCOCO [54], 3DPW [55], and UBody [53].
-Further qualitative comparisons are available in the Appendix.
-Methods AMASS [52] HPS [56]
-VPoser [35] 23.42 22.78
-Pose-NDF [36] 22.13 23.82
-MVAE [40] 26.80 N/A
-HuMoR [39] 22.69 N/A
-DPoser (ours) 19.87 20.54
-Table 4: Performance metrics (MPJPE) for motion denoising.
-5.4 Pose Completion
-Figure 6: Qualitative evaluations of pose completion. Three hypotheses are drawn from VPoser [35] and our DPoser.
-Timestep scheduling HMR Pose Completion Motion Denoising
-PA-MPJPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
-Random 68.21 98.57 109.53 102.37 43.33 23.87
-Fixed 57.72 85.11 40.44 36.70 45.69 22.54
-Full annealed 61.15 85.27 42.97 39.42 39.72 20.80
-Truncated annealed 57.28 84.79 39.77 36.15 38.21 19.87
-Table 5: Comparative performance of different timestep scheduling across pose-related tasks.
-Noise std AMASS [52] HPS [56]
-20.00 31.93/13.64 31.93/13.45
-40.00 63.81/19.87 63.81/20.54
-100.00 159.78/33.18 159.78/35.32
-Table 6: DPoser in motion denoising under varying noise scales. MPJPE is reported as before/after applying DPoser denoising.
-In practical scenarios like those encountered in the UBody dataset [53] (refer to Fig.
-4(d)), HMR algorithms often grapple with occlusions leading to incomplete 3D pose estimates.
-In this context, our ambition is to recover full 3D poses from such partially observed data, where the unseen parts are initialized with random noise.
-Given that Pose-NDF [36] and VPoser [35] have not been previously scrutinized for this task, we introduce two experimental paradigms—generation-based and optimization-based.
-In the optimization regimen, we incorporate an L2-norm fidelity term for the observed pose and constrain the entire pose using these deep priors.
-In contrast, the generation-based scheme for VPoser involves encoding the whole pose, followed by decoding and sampling from the resulting Gaussian distributions.
-For Pose-NDF, we iteratively project the whole pose onto the learned manifold, replacing the unmasked components with known data.
-For DPoser, we report optimization-based outcomes while the generation-based results employing other inverse problem solvers [8, 20, 19] can be checked in the Appendix.
-Given the task’s inherent ambiguity, we sample multiple solutions and select the best-fitting one, except the deterministic Pose-NDF.
+Moreover, GAN-S [9] implementations require a GAN-inversion phase to convert initial poses into their latent representations, which is notably time-consuming.
 Tab.
 2 and Fig.
-6 demonstrate DPoser’s superiority in pose completion under occlusions, offering multiple plausible solutions, unlike VPoser [35] and Pose-NDF [36], with the latter facing generalization issues when applied to such unseen noisy poses.
+5 showcase the comparative performance of DPoser, highlighting its exceptional ability in HMR tasks.
+Notably, when fitting from scratch, it surpasses established SOTA priors like GAN-S [9] and Pose-NDF [49] and rivals the specific regression-based model [25].
+The integration of CLIFF as initialization further amplifies DPoser’s performance, underscoring its efficiency and the benefits of employing refined starting conditions.
+Fig.
+6 further confirms DPoser’s superior efficacy and adaptability across multiple datasets including EHF [37], MSCOCO [27], 3DPW [51], and UBody [26].
+4.4 Pose Completion
+In practical scenarios like those encountered in the UBody dataset [26] (refer to Fig.
+5(d)), HMR algorithms often grapple with occlusions leading to incomplete 3D pose estimates.
+In this context, our ambition is to recover full 3D poses from partially observed data, initializing the occluded parts with random noise.
+Our DPoser model is employed to refine these initially implausible poses into feasible ones, utilizing an L2 loss on the visible parts to ensure data consistency.
+Initialization VPoser Pose-NDF DPoser
+Zeros 180.90 157.50 73.92
+10mm noise 181.86 172.50 74.69
+100mm noise 180.25 511.51 74.19
+Table 3: Pose completion on the AMASS [32] dataset (left leg under occlusion, single-hypotheses) using various initialization strategies. DPoser demonstrates its effectiveness across all conditions.
+In parallel, we employ a comparable optimization strategy for both Pose-NDF [49] and VPoser [37].
+Notably, Tab.
+3 reveals that Pose-NDF struggles with poorly initialized poses unseen during its training phase.
+To mitigate this issue, we have to initialize the occluded poses near zero (close to rest pose) for Pose-NDF to prevent optimization divergence.
+Additionally, as a task-specific baseline, we adapt the original VPoser model into CVPoser by incorporating conditional inputs within its VAE framework [22].
+This modification enables the encoder and decoder to process additional partial poses, facilitating end-to-end conditional sampling.
+Methods Occ. left leg Occ. legs Occ. arms Occ. trunk
+PoseNDF ($S=1$[49] 158.21 159.19 201.00 75.42
+PoseNDF ($S=5$) 147.66/158.11/7.62 151.86/159.21/5.33 196.36/200.92/3.30 70.88/75.39/3.25
+PoseNDF ($S=10$) 144.38/158.06/8.31 149.38/159.14/5.90 194.79/200.87/3.63 69.45/75.38/3.54
+VPoser (S=1) [37] 180.78 198.18 159.86 37.75
+VPoser ($S=5$) 167.92/181.30/10.53 178.77/198.15/14.51 148.17/159.65/8.64 31.83/37.79/4.54
+VPoser ($S=10$) 162.82/181.09/12.21 172.83/198.31/16.30 144.53/159.80/9.69 30.06/37.78/4.99
+CVPoser ($S=10$) ${}^{\dagger}$ 71.66/145.52/51.68 90.49/148.30/38.46 83.02/136.82/36.47 18.77/37.83/13.12
+DPoser(ours) ($S=1$) 74.48 97.39 81.49 28.58
+DPoser(ours) ($S=5$) 42.64/73.85/24.36 67.70/97.06/22.29 58.52/82.37/18.33 17.11/28.59/8.92
+DPoser(ours) ($S=10$) 35.37/74.01/26.47 59.25/96.77/24.55 51.27/81.76/20.04 13.95/28.57/9.85
+Table 4: Performance metrics (min/mean/std of MPJPE across multiple hypotheses) for pose completion under varying occlusion scenarios.
+$S$ denotes the number of hypotheses.
+${}^{\dagger}$ Task-specific baseline trained with partial poses as conditional input.
+Figure 7: Qualitative evaluations of pose completion.
+Three hypotheses are drawn for each method.
+DPoser uniquely offers multiple plausible solutions for partially observed poses, a scenario where competitors often struggle due to limited generalization.
+Given the inherent uncertainties within this task, we generate multiple solutions and evaluate them based on their minimum, mean, and standard deviation errors against the ground truth.
+As illustrated in Tab.
+4, DPoser exhibits superior performance across different occlusion scenarios compared to existing pose priors and even the task-specific CVPoser, highlighting its effectiveness in pose completion.
+The qualitative evaluations are presented in Fig.
+7.
+Here, we observe that DPoser can generate a multitude of plausible poses, a capability lacking in VPoser [37].
+Pose-NDF [49], meanwhile, struggles with generalizing to unseen noisy poses and making plausible adjustments from its rest pose initialization.
 4.5 Motion Denoising
@@
 The task aims to estimate clean body poses from noisy 3D joint positions in motion capture sequences.
-Adhering to the setup outlined in HuMoR [39], we utilize 60-frame sequences from the AMASS [52] test dataset and artificially introduce Gaussian noise with a standard deviation of 40 mm to the 3D joint positions.
+Adhering to the setup outlined in HuMoR [40], we utilize 60-frame sequences from the AMASS [32] dataset and artificially introduce Gaussian noise with a standard deviation of 40 mm to the 3D joint positions.
 Moreover, we conduct experiments on HPS datasets [15] without additional training to validate the generalization.
-Our optimization loss is composed of a data fidelity term, a temporal consistency term, and our pose prior, as discussed in the Appendix.
 As presented in Tab.
-4, DPoser sets a new standard in motion denoising, outperforming even specialized motion priors like HuMoR [39].
+5, DPoser sets a new standard in motion denoising, outperforming even specialized motion priors like HuMoR [40].
 To further confirm the robustness of DPoser, we conduct evaluations under varying conditions to gauge DPoser’s denoising capabilities.
 The results, detailed in Tab.
-6, reveal that DPoser consistently achieves significant reductions in MPJPE, maintaining robust performance even under extreme noise conditions, specifically at a noise standard deviation of 100.00 mm.
-We direct the reader to our Appendix for qualitative assessments.
+6, reveal that DPoser consistently achieves significant reductions in MPJPE, maintaining robust performance under extreme noise conditions.
 4.6 Ablation Study
+Methods AMASS [32] HPS [15] No prior 24.19 23.67 VPoser [37] 23.42 22.78 Pose-NDF [49] 22.13 21.60 MVAE [28] 26.80 N/A HuMoR [40] 22.69 N/A DPoser (ours) 19.87 20.54 Table 5: Performance metrics (MPJPE) for motion denoising under 40 mm noise.
+Noise std AMASS [32] HPS [15] 20.00 31.93/13.64 31.93/13.45 40.00 63.81/19.87 63.81/20.54 100.00 159.78/33.18 159.78/35.32 Table 6: DPoser in motion denoising under varying noise scales.
+MPJPE is reported as before/after applying DPoser denoising.
+Timestep scheduling HMR Pose Completion Motion Denoising
+PA-MPJPE $\downarrow$ MPJPE ($S=10$) $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
+Random 58.84 86.23/121.57/23.16 43.33 23.87
+Fixed 56.55 36.99/71.68/23.41 45.69 22.54
+Uniform 59.28 42.72/75.70/21.84 39.72 20.80
+Truncated 56.05 35.37/74.01/26.47 38.21 19.87
+Table 7: Evaluation of timestep scheduling strategies on key pose-related tasks, highlighting the superior efficacy of the proposed truncated scheduling.
 In our ablation study, we initially focus on the impact of truncated timestep scheduling on DPoser’s performance.
-This involves contrasting our proposed scheduling strategy against three established methods—random, fixed, and full annealed scheduling [62, 18, 19, 8].
+This involves contrasting our proposed scheduling strategy against three established methods—random, fixed, and uniform scheduling [34, 33, 6, 48].
 As Tab.
 7 demonstrates, our strategy consistently outperforms these alternatives across all evaluated tasks.
-Further, we delve into the training aspects of DPoser, such as rotation representations and the integration of an auxiliary loss akin to HuMoR [39].
-Additionally, using the same trained prior, we benchmark DPoser against SOTA diffusion-based inverse problem solvers [8, 20, 19] on the pose completion task.
-The results underscore that DPoser not only proves versatile across multiple pose-related tasks but also excels in performance.
+Additionally, we delve into the training aspects of DPoser, such as rotation representations and the integration of an auxiliary loss akin to HuMoR [40].
+Using the same trained prior, we also compare DPoser’s capabilities with SOTA diffusion-based solvers [48, 7, 6] on pose completion, revealing its superior versatility and performance.
 Detailed findings and analyses from these ablation studies are presented in the Appendix.
 5 Conclusion
-We introduce DPoser, a robust and vetsatile human pose prior framework grounded in diffusion models.
-Engineered for flexibility, DPoser shines across multiple pose-related tasks, enhanced by our innovative truncated timestep scheduling for test-time optimization.
-Comprehensive experiments substantiate DPoser’s superior performance over existing state-of-the-art pose priors, establishing it as an effective solution for a wide range of applications.
+We introduce DPoser, to our best knowledge, the first unconditional diffusion-based pose prior, tailored for an expansive array of pose-related tasks.
+Engineered for flexibility, DPoser can be implemented as a straightforward L2-loss regularizer and enhanced by our innovative truncated timestep scheduling for test-time optimization.
+Comprehensive experiments substantiate DPoser’s superior performance over existing state-of-the-art pose priors.
 Limitation and future work.
@@
 Future research could look into enhancing solution diversity via techniques like particle-based variational inference [30, 53].
-Furthermore, within the broader context of inverse problems, a plethora of existing methods [22, 23, 25, 24] could be adapted to leverage our diffusion-based prior.
+Furthermore, within the broader context of inverse problems we have framed, a plethora of existing methods [45, 2, 5, 35] could be adapted to leverage our diffusion-based prior.
 Exploring these methods holds great potential for future progress.
-Supplementary Material:
-DPoser: Diffusion Model as Robust 3D Human Pose Prior
+Ethical Considerations. For a discussion on the potential negative impacts of our work, please refer to the Appendix.
+Appendix for DPoser: Diffusion Model as
+Robust 3D Human Pose Prior
+In this appendix, we first briefly recap the parameterization of diffusion models and their connection to score functions in Sec.
+A, followed by the perspective of Score Distillation Sampling (SDS) to understand our DPoser regularization in Sec.
+B.
+We detail the experimental setup and nuances in Sec.
+C and dissect various training aspects of DPoser in Sec.
+D.
+The exploration of extended optimization techniques is discussed in Sec.
+E, and considerations for truncated timestep scheduling in image domains are presented in Sec.
+F.
+Additional qualitative results are showcased in Sec.
+G.
+Lastly, potential negative impacts such as biases in data and ethical concerns in application are considered in Sec.
+H.
 A Parameterization of Score-based Diffusion Models
@@
 (16) aligns precisely with the noise prediction form of diffusion models [16] (refer to Eq.
-(4)).
+(4) in the main text).
 This implies that by training $\epsilon_{\theta}(\mathbf{x}_{t};t)$ in a diffusion model context, we simultaneously get a handle on the score function, approximated as $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\approx-\frac{\epsilon_{\theta}(\mathbf{x}_{t};t)}{\sigma_{t}}$.
@@
 Strategy HMR Pose Completion Motion Denoising
-PA-MPJPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
-1step 57.28 84.79 39.77 36.15 38.21 19.87
-5steps 57.61 85.14 41.30 37.58 40.22 21.21
-10steps 57.66 85.07 41.28 37.39 40.69 21.34
+PA-MPJPE $\downarrow$ MPJPE ($S=10$) $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
+1 step 56.05 35.37/74.01/26.47 38.21 19.87
+5 steps 56.16 36.59/80.82/31.22 40.22 21.21
+10 steps 56.18 36.78/82.59/32.32 40.69 21.34
 Table S-1: Efficacy of different denoising steps in DPoser’s optimization.
-Interestingly, the gradient of DPoser in Eq. (10) coincides with Score Distillation Sampling (SDS) [27, 28], which can be interpreted as aiming to minimize the following KL divergence:
+Interestingly, the gradient of DPoser (Eq. (10) in the main text) coincides with Score Distillation Sampling (SDS) [38, 52], which can be interpreted as aiming to minimize the following KL divergence:
 $KL\big{(}p_{0t}\left(\mathbf{x}_{t}\mid\mathbf{x}_{0}\right)\parallel p_{t}^{\mathtt{SDE}}\left(\mathbf{x}_{t};\theta\right)\big{)},$ (17)
@@
 A plausible explanation could be that our proposed truncated timestep scheduling effectively manages low noise levels (i.e., small $t$), thus negating the need for more denoising steps.
+In addition, iterative denoising in each optimization step may cause error accumulations, leading to inaccurate gradients.
 C Experimental Details
-This section elaborates on the specifics of our pose completion and motion denoising experiments, and also highlights DPoser’s proficiency in pose interpolation tasks.
+This section elaborates on the specifics of our pose completion and motion denoising experiments.
 C.1 Pose Completion
 For partial observations $\mathbf{y}$, the measurement operator $\mathcal{A}$ is modeled as a mask matrix $M\in\mathbb{R}^{d\times n}$.
-Based on our optimization framework (i.e., Algorithm 1), we define the task-specific loss, $L_{\text{comp}}$, as follows:
+Based on our optimization framework (Algorithm 1 in the main text), we define the task-specific loss, $L_{\text{comp}}$, as follows:
 $L_{\text{comp}}=||M\mathbf{x}_{0}-\mathbf{y}||_{2}^{2}.$ (18)
-Here, $\mathbf{x}_{0}$ denotes the whole body pose $\theta$ we try to recover, where the unseen parts are initialized as random noise.
-In ourthe following ablated studies, if not specified, the evaluation is performed using 10 hypotheses on the AMASS [52] test set with left leg occlusion.
-C.2 Motion denoising
+Here, $\mathbf{x}_{0}$ denotes the complete body pose $\theta$ we try to recover, where the unseen parts are initialized as random noise.
+In the following ablated studies, if not specified, the evaluation is performed using 10 hypotheses on the AMASS [32] dataset with left leg occlusion.
+C.2 Motion Denoising (Noisy Input)
 Methods AMASS [32] HPS [15]
 20mm 100mm 20mm 100mm
+No prior 15.33 51.48 16.26 50.87
 VPoser [37] 15.20 49.10 17.24 46.69
@@
 where $M_{J}$ denotes the 3D joint positions regressed from SMPL [31] and $\beta_{0}$ is the constant mean shape parameters.
-In complement to the comparative analysis presented in Tab.
-4, we extend our evaluation to include scenarios with varying noise levels.
+In complement to the comparative analysis presented in Table 4 of our main text, we extend our evaluation to include scenarios with varying noise levels.
 This extended examination, detailed in Tab.
 S-2, showcases DPoser’s exceptional performance against state-of-the-art (SOTA) pose priors, especially under conditions of high noise, manifesting DPoser’s resilience to noise.
-C.3 Pose interpolation
-(a) Pose reconstruction.
-(b) Interpolation between two poses.
-Figure S-1: Visual results of pose reconstruction and interpolation.
-For the forward SDE (Eq. (1)), there exists a probability flow ODE with the same marginal probability densities corresponding to it, expressed as: [8]
-$\mathrm{d}\mathbf{x}=[\mu(t)\mathbf{x}-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}\left(\mathbf{x}_{t}\right)]\mathrm{d}t.$ (21)
-By replacing $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ with our approximation $-\frac{\epsilon_{\theta}(\mathbf{x}_{t};t)}{\sigma_{t}}$, this ODE can be treated as a neural ODE [5], which is fully invertible.
-By integrating through it, we can encode any data point $\mathbf{x}_{0}$ into its latent representation $\bar{\mathbf{x}}_{1}$ and decode it through its reverse-time counterpart.
-For interpolation, we start with two poses, $\mathbf{x}_{0}^{0}$ and $\mathbf{x}_{0}^{1}$, and encode them into latent representations $\bar{\mathbf{x}}_{1}^{0}$ and $\bar{\mathbf{x}}_{1}^{1}$.
-Utilizing spherical linear interpolation (Slerp), we obtain interpolated latents $\bar{\mathbf{x}}_{1}^{t}=\text{Slerp}(\bar{\mathbf{x}}_{1}^{0},\bar{\mathbf{x}}_{1}^{1};t)$ for $t\in[0,1]$, which are then decoded to produce novel poses.
-The Slerp function is defined as:
-$\text{Slerp}(A,B;t)=\frac{\sin((1-t)\omega)}{\sin\omega}A+\frac{\sin(t\omega)}{\sin\omega}B,$ (22)
-where $\omega=\arccos(A\cdot B/\|A\|\|B\|)$.
-To validate the informativeness of our encoding, we employ MPJPE and MPVPE metrics for reconstruction evaluation.
-Compared to VPoser [35], a VAE-based prior, DPoser exhibits superior performance with 4.79mm and 6.58mm errors, against 9.85mm and 12.16mm, respectively.
-Qualitative results further confirm DPoser’s strong reconstruction and smooth interpolation capabilities, as shown in Fig.
-S-1.
-D Ablated DPoser Training
+C.3 Motion Denoising (Partial Input)
+This task focuses on reconstructing clean poses, $\theta^{t}$, from partially observed joint positions, $J_{\text{obs}}^{t}$, across $N$ frames, employing a known mask matrix to identify visible joints.
+The optimization objective mirrors that of motion denoising (Sec.
+C.2), but incorporates a mask in Eq.
+(19) to specifically target visible parts, ensuring that only these segments guide the recovery process.
+We conducted experiments on the AMASS dataset [32] to assess our model’s performance on this task with two types of occlusions: legs and left arm.
+The quantitative results of these experiments are detailed in Tab.
+S-3, and the accompanying visualizations are provided in Sec.
+G.
+In leg occlusion scenarios, the AMASS dataset primarily showcases straight poses, offering minimal diversity.
+This scenario permits decent outcomes without incorporating a pose prior, since the optimization’s starting point closely aligns with these prevalent poses.
+However, VPoser’s mean-centered characteristic hinders its ability to faithfully replicate the visible areas.
+On the other hand, Pose-NDF falls short in enhancing the occluded parts.
+DPoser accurately handles visible parts and guides occluded ones for more realistic poses.
+For left arm occlusions, which involve more varied movements, DPoser markedly surpasses other methods, underlining its adaptability and precision in handling diverse motion patterns.
+Methods Occlusion MPJPE MPVPE
+Vis. Occ. All. All.
+No prior Legs 0.26 14.72 5.52 5.45
+VPoser Legs 1.75 14.29 6.31 7.38
+PoseNDF Legs 0.25 15.71 5.87 5.64
+DPoser (ours) Legs 0.28 12.24 4.63 3.65
+No prior Left Arm 0.26 24.87 4.74 9.91
+VPoser Left Arm 1.21 13.23 3.40 7.68
+PoseNDF Left Arm 0.25 17.70 3.42 7.86
+DPoser (ours) Left Arm 0.27 7.80 1.64 3.81
+Table S-3: Comparative analysis of methods for motion denoising with different occlusions (Legs and Left Arm) on the AMASS dataset.
+Errors (in cm) are evaluated in terms of MPJPE across visible (Vis.), occluded (Occ.), and all joints, along with MPVPE for all vertices.
+D Ablated DPoser’s Training
 Normalization HMR Pose Completion Motion Denoising
-PA-MPJPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
-w/o norm 58.62 83.75 60.70 55.47 44.82 24.04
-min-max 59.55 85.63 58.18 53.06 42.70 21.29
-z-score 57.66 85.47 52.86 48.27 38.57 20.24
-Table S-3: Evaluation of DPoser’s performance under different normalization methods, specifically for the axis-angle rotation representation.
+PA-MPJPE $\downarrow$ MPJPE ($S=10$) $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
+w/o norm 57.88 45.37/102.28/41.08 44.82 24.04
+min-max 59.17 47.41/107.00/43.42 42.70 21.29
+z-score 56.49 34.37/72.47/26.32 38.57 20.24
+Table S-4: Evaluation of DPoser’s performance under different normalization methods, specifically for the axis-angle rotation representation.
 Representation HMR Pose Completion Motion Denoising
-PA-MPJPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
-axis-angle 57.28 84.79 52.14 47.60 38.21 19.87
-6D rotations 58.02 85.93 56.28 51.54 38.44 20.12
-Table S-4: Comparative performance of rotation representations under z-score normalization across multiple tasks and metrics.
-Figure S-2: Evolution of MPJPE during DPoser training, evaluated on pose completion.
-The evaluation is conducted on the AMASS [52] test set, utilizing 10 hypotheses and simulating occlusion scenarios affecting both legs.
-Building on the insights shared in Sec.
-4.1 in our main text, this part dissects the impact of different rotation representations and normalization techniques on DPoser’s performance.
+PA-MPJPE $\downarrow$ MPJPE ($S=10$) $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
+axis-angle 56.05 34.76/72.41/26.09 38.21 19.87
+6D rotations 57.54 40.89/81.43/27.31 38.44 20.12
+Table S-5: Comparative performance of rotation representations under z-score normalization across multiple tasks and metrics.
+This section dissects the impact of different rotation representations and normalization techniques on DPoser’s performance.
 Initially, we examine axis-angle representation, comparing various normalization strategies: min-max scaling, z-score normalization, and no normalization.
 Our findings, summarized in Tab.
-S-3, indicate that z-score normalization is generally the most effective.
+S-4, indicate that z-score normalization is generally the most effective.
 Subsequently, using this optimal normalization, we explore 6D rotations [57] as an alternative.
 As evidenced by Tab.
-S-4, axis-angle representation offers superior performance.
-This preference can be attributed to the effective modeling capabilities of diffusion models, along with the inherent advantages of axis-angle in capturing bounded joint rotations.
-Inspired by HuMoR [39], we experiment with integrating the SMPL body model [33] as a regularization term during DPoser’s training.
-Alongside the prediction of additive noise, as outlined in Eq (4), we employ a 10-step DDIM sampler [11] to recover a “clean” version of the pose, denoted as $\tilde{\mathbf{x}}_{0}$, from the diffused $\mathbf{x}_{t}$.
+S-5, axis-angle representation offers superior performance.
+This preference can be attributed to the effective modeling capabilities of diffusion models, along with the inherent advantages of axis-angle in capturing bounded joint rotations for regression tasks like human mesh recovery.
+Inspired by HuMoR [40], we experiment with integrating the SMPL body model [31] as a regularization term during training.
+Alongside the prediction of additive noise, as outlined in Equation (4) in the main text, we employ a 10-step DDIM sampler [44] to recover a “clean” version of the pose, denoted as $\tilde{\mathbf{x}}_{0}$, from the diffused $\mathbf{x}_{t}$.
 The regularization loss aims to minimize the discrepancy between the original and recovered poses under the SMPL body model $M$:
-$L_{\mathrm{reg}}=$ $||M_{J}(\tilde{\mathbf{x}}_{0},\beta_{0})-M_{J}(\mathbf{x}_{0},\beta_{0})||_{2}^{2}$
-$+||M_{V}(\tilde{\mathbf{x}}_{0},\beta_{0})-M_{V}(\mathbf{x}_{0},\beta_{0})||_{2}^{2}.$
+$L_{\mathrm{reg}}=||M_{J}(\tilde{\mathbf{x}}_{0},\beta_{0})-M_{J}(\mathbf{x}_{0},\beta_{0})||_{2}^{2}+||M_{V}(\tilde{\mathbf{x}}_{0},\beta_{0})-M_{V}(\mathbf{x}_{0},\beta_{0})||_{2}^{2}.$ (21)
 Here, $\beta_{0}$ represents the mean shape parameters in SMPL.
 To account for denoising errors, we scale the regularization loss by $\mathrm{log}(1+\frac{\alpha_{t}}{\sigma_{t}})$, thereby increasing the weight for samples with smaller $t$ values (less noise).
-Fig.
-S-2 visualizes the impact of this regularization on MPJPE during the training, specifically for pose completion tasks with occlusion of both legs.
+Fig. S-1 visualizes the impact of this regularization on MPJPE during the training, specifically for pose completion tasks with occlusion of both legs.
+Figure S-1: MPJPE evolution in DPoser training for pose completion, assessed on AMASS [32] with 10 hypotheses under legs occlusion scenarios.
 We observe that weighted regularization offers slight performance gains in the early training process, while the absence of weighting introduces instability and deterioration in results.
@@
 Therefore, we opted not to include this regularization in our main experiments.
-E Extended DPoser Optimization
+E Extended DPoser’s Optimization
 Methods Occ. left leg Occ. legs Occ. arms Occ. trunk
-MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$ MPVPE $\downarrow$ MPJPE $\downarrow$
-ScoreSDE [8] 52.86 48.27 80.68 74.74 82.38 65.81 27.97 17.27
-DPS [19] 41.99 38.08 66.16 61.28 64.75 51.75 25.68 15.90
-MCG [20] 51.97 47.46 80.07 74.14 82.22 65.58 27.93 17.27
-DPoser (ours) 39.77 36.15 65.24 60.54 66.77 52.75 21.58 13.41
-Table S-5: Comparative evaluation of diffusion-based solvers for pose completion.
+ScoreSDE [48] 48.73/106.32/41.30 74.68/128.32/37.27 66.89/127.86/48.15 16.69/34.54/12.21
+DPS [6] 40.51/104.32/54.57 64.26/113.46/33.71 60.63/119.85/42.78 15.10/33.90/13.27
+MCG [7] 49.04/106.37/41.07 74.90/128.53/37.40 66.17/127.72/48.15 16.69/34.66/12.23
+DPoser(ours) 35.37/74.01/26.47 59.25/96.77/24.55 51.27/81.76/20.04 13.95/28.57/9.85
+Table S-6: Comparative evaluation of diffusion-based solvers for pose completion on the AMASS dataset [32] (hypotheses number $S=10$).
 In addressing pose-centric tasks as inverse problems, we propose a versatile optimization framework, which employs variational diffusion sampling as its foundational approach [33].
@@
 Our findings, captured in Tab.
-S-5, reveal that DPoser outperforms the competitors under most occlusion conditions.
+S-6, reveal that DPoser outperforms the competitors under most occlusion conditions.
 Consequently, DPoser emerges not merely as a universally applicable solution to pose-related tasks, but also as an exceptionally efficient one.
@@
 Although some state-of-the-art techniques [5, 35] offer solutions by jointly modeling operator $\mathcal{A}$ and data distributions, a full-fledged discussion on this subject is beyond this paper’s purview and remains an open question for future work.
-F More Qualitative Results
-We show more qualitative results for pose generation (Fig. S-3), pose completion (Fig. S-4), human mesh recovery (Fig. S-5), and motion denoising (Fig. S-6).
+F Truncated Timestep Scheduling on Images
+Figure S-2: Image inpainting using standard (a) and truncated (b) timestep scheduling.
+The process evolution is shown over iterations with the middle row depicting the log-magnitude spectrum and the bottom row the phase spectrum.
+Exploring truncated timestep scheduling for image-based tasks, we find its suitability for human poses doesn’t translate well to images.
+Initial timesteps are critical in image domains for generating foundational perceptual content.
+In our study, we employed a 256x256 unconditional diffusion model [11] trained on ImageNet [10] with variational diffusion sampling [33] for image inpainting.
+Comparing standard (timesteps 990 to 0) and truncated scheduling (timesteps 495 to 0), both with 100 steps, the experiments confirmed that truncation compromises image quality (Fig.
+S-2).
+The standard approach preserved perceptual content, while truncation produced disjointed patches, misaligned with the original image context.
+These results affirm that truncated timestep scheduling excels in pose data where key information emerges in later stages but falls short in image tasks where early timesteps are essential.
+This scheduling is thus bespoke to the characteristics of human pose estimation and is unsuitable for image processes that rely on the full diffusion timeline for content fidelity.
+G More Qualitative Results
+We show more qualitative results for pose generation (Fig. S-3), pose completion (Fig. S-4), human mesh recovery (Fig. S-5) and motion denoising (Fig. S-6, Fig. S-7).
+H Potential Negative Impacts
+  • Bias and Fairness Concerns: Human pose prior learning models may inadvertently encode biases present in the training data, leading to biased predictions or discriminatory outcomes.
+This can perpetuate existing societal biases and inequalities, particularly if the training data is not representative or balanced across diverse demographics.
+  • Ethical Considerations: The use of human pose prior learning models in applications such as surveillance, security, or healthcare raises ethical concerns regarding individual privacy, autonomy, and consent.
+There are debates about the appropriate use of such technologies and the potential for unintended consequences or misuse.
+  • Dependency on Data Quality: Human pose prior learning models heavily rely on the quality and diversity of the training data.
+Poorly annotated or biased datasets can negatively impact the performance and reliability of these models, leading to inaccurate or unreliable predictions.
 Figure S-3: Pose generation. DPoser can generate diverse and realistic poses.
-Figure S-4: Pose completion. (a) Right arm under occlusion. (b) Trunk under occlusion.
+Figure S-4: Pose completion. (a) Left leg under occlusion. (b) Trunk under occlusion.
 Figure S-5: Human mesh recovery.
@@
 […]
+[…]